### v2

# Measuring the (Un)Faithfulness of Concept-Based Explanations

Abstract
Post-hoc, unsupervised concept-based explanation methods (U-CBEMs) translate a vision model’s internal reasoning into human-understandable concepts, leading to interpretable explanations. However, we find that many state-of-the-art (SOTA) U-CBEMs are not faithful: their concepts seem interpretable but fail to reproduce the model’s predictions. We argue that this deficiency has gone unnoticed due to fragmented evaluation—each paper proposes its own faithfulness measure, with no measure-over-measure comparison or broad benchmarking. We close this gap by (i) organizing prior metrics in a unified framework, discussing their limitations, and identifying desiderata for a faithfulness measure; (ii) introducing the Surrogate Faithfulness (SURF) measure, which quantifies faithfulness via the predictive loss of a surrogate that maps explanations to the model’s outputs; and (iii) delivering the first comprehensive U-CBEM faithfulness benchmark across diverse tasks and architectures. In a controlled setting, SURF outperforms prior faithfulness measures in measure-over-measure comparisons, and applying SURF to SOTA U-CBEMs reveals that many visually appealing U-CBEMs are surprisingly unfaithful. We demonstrate SURF applicability in two downstream settings—(i) faithfulness versus the number of concepts used in the explanation and (ii) U-CBEM robustness to adversarial attacks—underscoring SURF’s value as a reliable faithfulness measure. Code to be released.
1 Introduction
Deep learning models have delivered state-of-the-art results across diverse tasks, yet their internal reasoning remain difficult to interpret [12], which is especially problematic in high-stakes domains such as healthcare and finance. This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior. Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that accurately reflect the model’s internal reasoning and allow humans to reliably predict the model’s output. XAI methods must balance the natural tension between interpretability and faithfulness. While interpretability is typically assessed with human studies, faithfulness can be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model. Because explanations are inevitably lossy, some discrepancy between the surrogate and the model is expected; a faithfulness measure must therefore specify both (i) an appropriate surrogate and (ii) a metric for the induced loss.
One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model. Since the explanation and the model are merged, these approaches eliminate the need for faithfulness checks; however, they often underperform black-box models by reducing model complexity (capacity) for explainability. In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution—commonly found through gradients or perturbations [38, 30, 23, 28, 7]. For images, these attributions are visualized as pixel-level heatmaps. Although such explanations are intuitive, attributions are usually unfaithful, since they are estimated from noisy gradients in deep neural networks or from unrealistic perturbations on the input space [1, 11, 32, 15]. The heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [26, 4].
Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [10, 3, 17, 46]. To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss aspects of the model’s reasoning, raising faithfulness concerns [29]. To avoid annotation, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output.
We argue that, in practice, U-CBEMs have prioritized interpretability at the expense of faithfulness, yielding explanations that look compelling but are surprisingly unfaithful. Despite the importance of faithful explanations, we find that all prior works introduce new faithfulness measures, with no measure-over-measure comparison and sparse cross-U-CBEM evaluations. Thus, there is a lack of consensus on faithfulness measures for U-CBEMs, limiting our understanding of how faithful current approaches are.
This work serves to address this gap. Our contributions are to:
  1. 1.
    Organize prior U-CBEM faithfulness measures under a common framework, allowing us to discuss limitations in current measures and identify appropriate desidarata for a faithfulness measure.
  2. 2.
    Propose Surrogate Faithfulness (SURF), a faithfulness measure satisfying the desidarata. We conduct a measure-over-measure comparison to empirically validate our faithfulness measure against prior measures.
  3. 3.
    Conduct the first, comprehensive faithfulness benchmark of current U-CBEMs across a variety of tasks and model architectures. Most SOTA U-CBEMs that were thought to be faithful are found not to be.
  4. 4.
    Extend SURF to two downstream applications. First, studying faithfulness as a function of the number of concepts shows that some U-CBEMs do not get more faithful as they discover more concepts. Second, analyzing the faithfulness of U-CBEMs when the model is under adversarial attack reveals robustness issues.
2 Preliminaries
We first introduce notation that will be used in the remainder of the paper, as well as preliminaries to familiarize the reader with the field.
Background: Model $\phi:\mathcal{X}\to\mathcal{Y}$ maps from an input space $\mathcal{X}\subseteq\mathbb{R}^{P}$ to an output space $\mathcal{Y}\subseteq\mathbb{R}^{C}$. Assume $\phi$ admits an intermediate space $\mathcal{H}\subseteq\mathbb{R}^{D}$. Let $g:\mathcal{X}\to\mathcal{H}$ and $f:\mathcal{H}\to\mathcal{Y}$. Thus, $\phi(\mathbf{X})=f(g(\mathbf{X}))$. Let $g(\mathbf{x})=\mathbf{h}\in\mathcal{H}$ represent the embedding of $\mathbf{x}$ (e.g., image patch), and $g(\mathbf{X})=\mathbf{H}\in\mathbb{R}^{(HW)\times D}$ denotes the vectorized application of $g$ on $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{HW}\end{bmatrix}^{T}$, where $H,W$ are the spatial dimensions.
Let $E$ be an explanation function, and let $\mathbf{e}=E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image). Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate. Motivated by Ribeiro et al. [30], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$. Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
$Faith(E;d,s)=\int_{\mathcal{X}}d(\phi(\mathbf{X}),s(E(\mathbf{X})))d\mathbf{X}$ (1)
U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and their concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$. Note that while $K$ can vary per output, all prior works find the same number of CAVs for each output. U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$. The explanation for output $i$ is:
$E_{i}(\mathbf{X})$ $=E_{i}(\mathbf{X};g,V_{i},A_{i},\mathcal{P})$ (2)
$=\{\mathcal{P}(g(\mathbf{X});V_{i}),A_{i}\}$
$E_{i}(\mathbf{X})$ is then conveyed to the user through some visual interface. Note that $\mathcal{P}(g(\mathbf{X}))$ denotes the vectorized application of $\mathcal{P}$ on $g(\mathbf{X})=\mathbf{H}$.
Following prior works, we study faithfulness in the final layer, where the spatial dimensions $H=W=1$, so $\mathbf{H}=\mathbf{h}$. Given that U-CBEMs operate on $g(X)$, the faithfulness measure can be simplified to:
$Faith_{\text{U-CBEM}}(\mathcal{P},\{V_{i}\}_{i=1}^{C},\{A_{i}\}_{i=1}^{C};d,s)=$ (3)
$\int_{\mathcal{H}}d(f(\mathbf{h}),\{s(\mathcal{P}(\mathbf{h};V_{i}),A_{i})\}_{i=1}^{C})d\mathbf{h}$
3 Related Works
3.1 Unsupervised Concept-Based Explanations
Wang et al. [37] initiated this line of work, finding CAVs by K-Means clustering intermediate embeddings computed on a dataset. Each cluster was associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches. Ghorbani et al. [13] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate. To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al. [39] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [41] to find concept importance. ICE [44] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation. The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA). Building on this, Fel et al. [9] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance. CDISCO [14] uses the singular value decomposition (SVD) to find CAVs and a gradient-based method for finding concept importance. Recently, MCD [35] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
Table 1: Deletion-Based Measures: We organize deletion-based measures introduced by prior works according to their differing factors. The factors we use are Perturbation Space, Deletion Strategy, and how the results are presented (Y-axis and X-axis). Notice how all prior works introduce a different faithfulness measure, with no measure-over-measure comparison.
Method Perturbation Space Deletion Strategy Y-axis X-axis
ACE Pixel Constant (grey) Accuracy # Deleted Concepts
CDISCO (M1) Pixel Constant (grey) # Classes with $80\%$ accuracy loss # Deleted Concepts
MCD Pixel Inpainting Accuracy Fraction of Inpainted Pixels
CRAFT Concept Constant (zero) Change in Class Logit Score # Deleted Concepts
CDISCO (M2) Weight Constant (zero) Accuracy # Deleted Concepts
3.2 Concept Bottleneck Models
U-CBEMs are not to be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model. These methods [20, 18, 42, 27] train a model to first predict a set of interpretable concepts (e.g., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification). By doing so, CBMs fully decomposes the model’s prediction into interpretable concepts, and exposing the concept representation allows users to intervene (e.g., increase the strength of the “red” concept to help the model recognize the “stop sign”). While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts. As CBMs are inherently explainable, the notion of faithfulness does not extend to them. This work will only discuss faithfulness as it pertains to U-CBEMs.
4 Unifying U-CBEM Faithfulness Measures
We provide a comprehensive review of all faithfulness measures that were introduced by the works mentioned in Sec. 3.1. We find that each work uses a different measure for faithfulness, indicating a lack of unification in the field and preventing fair comparisons of faithfulness across different U-CBEMs. As part of our unified framework, we place prior faithfulness measures into two categories: 1) deletion-based and 2) surrogate-based. We detail the dimensions across which prior deletion-based measures differ; this is followed by a discussion on their limitations, which raise questions on the validity of past assessments. Then, we similarly characterize prior surrogate-based measures and demonstrate their drawbacks.
4.1 Deletion-Based Measures
Deletion-based measures evaluate faithfulness by observing the effect of deleting a concept on the model’s output. Intuitively, these measures assert that removing the most important concepts will result in the greatest model degradation. After obtaining CAVs and their importances (for a class), concepts are removed (Deletion Strategy) from the input’s Perturbation Space in decreasing order of importance, and the effect on the model is measured after each deletion. Results are visualized by plotting the effect on the model (Y-axis) with respect to some measure of how many concepts were deleted (X-axis). Prior works make different choices along each italicized dimension mentioned.
Perturbation Space: defines the space in which concepts are removed. This is either the image space, the model’s weight space, or the concept space. If making removals in the pixel or the weight space, the U-CBEM must have some mechanism to associate the concept to this space (e.g., which pixels in the image represent the concept).
Deletion Strategy: defines how concepts are removed. Concepts are usually removed from the Perturbation Space by setting the concept to a baseline value (e.g., 0).
Y-axis: defines what is measured after each deletion. Typically, this is accuracy (e.g., in object classification).
X-axis: represents how much “information” was deleted. This is typically the number of deleted concepts.
To understand the differences between prior deletion measures, Tab. 1 organizes prior measures according to these dimensions. ACE [13] deletes concepts in the image space by setting associated pixels to a constant value (Constant Deletion). They measure the model’s prediction accuracy after each deletion across a dataset. CRAFT [9] transforms the model’s embeddings into a concept representation. It makes Constant Deletions in the concept space by setting the concept to zero; then, they reconstruct the embedding from the perturbed concept representation. They measure the change in the corresponding class’s logit score. CDISCO [14] has two measures. The first measure (M1) makes Constant Deletions in the pixel space and measures the number of classes that have a greater than 80% accuracy degradation. The second measure (M2) deletes concepts by zeroing out parameters in the model’s weight space that are associated with the concept, and they plot classification accuracy. MCD [35] deletes concepts in the pixel space by using inpainting. They measure model classification accuracy with respect to the fraction of inpainted pixels, unlike all prior measures that use the number of deleted concepts.
4.1.1 Limitations of Deletion-Based Measures
There are two notable issues with deletion-based metrics that prevent them from accurately evaluating faithfulness. First, it is unclear how to delete a concept. Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0. The baseline must entirely delete the presence of the concept without affecting other concepts. However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [11]. Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
Second, there are no guarantees that representations in the Perturbation Space will stay on manifold after deletion. The feature attribution literature [11, 2, 34, 33, 40] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold. In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features. Thus, measuring feature importance with off-manifold inputs can lead to incorrect explanations [11]. More seriously, these explanations can be adversarially attacked; in [33], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs. Such issues can be mitigated by generating explanations using on-manifold input perturbations but requires accurate modeling of the input distribution [11, 34, 36, 43, 15].
Deletions in the concept or weight space manipulate intermediate representations, so the same concerns hold. Thus, we argue that deletion metrics have significant unresolved issues that introduce unreliability in their faithfulness evaluations, so we avoid using these approaches.
4.2 Surrogate-Based Measures
Surrogate-based measures evaluate faithfulness by approximating (Eq 3). Any surrogate-based measure must make two choices: 1) what the surrogate $s$ is, and 2) what the metric $d$ is. We will introduce prior faithfulness measures, and then we discuss their limitations.
4.2.1 Desidarata
The faithfulness definition in Eq. 3 does not constrain the surrogate $s$, but there are certain properties our surrogate should have. First, $s$ should be as simple as possible. Recognize that $s$ reflects the mental computation any human interpreter may be expected to do when trying to map the explanation onto the model’s prediction. An interpretable explanation that requires a complex $s$ is effectively useless; it will not give the human interpreter any meaningful insight into how the model reaches its prediction. Second, the $s$ should incorporate all components of the explanation. For example, U-CBEMs explain with both $V_{i}$ and $A_{i}$, so both should be used by the surrogate. If a component is not included, its impact on faithfulness cannot be measured. Third, we desire metric(s) $d$ that rewards explanations that can closely reconstruct all model outputs (e.g. across all classes), instead of only the model’s predicted class. This directly follows from Eq. 3.
Table 2: Surrogate-Based Measures: We compare prior surrogate-based measures introduced in ICE and C-SHAP to the proposed Surrogate Faithfulness (SURF) measure. According to our desidarata, we require surrogates that are simple (e.g., not reconstruction-based and not requiring additional parameters), incorporate concept importances into the surrogate, and measure errors across all outputs. Prior works fail to meet these desidarata (except for ICE not requiring additional parameters). We mark in green and red the properties that meet our desidarata. A ✓  or ✗  denotes that the method does or does not have the property.
Property ICE C-SHAP SURF (Ours)
Reconstruction-Based?
Additional Parameters?
Uses Concept Importance?
Errors on All Outputs?
4.2.2 Prior Surrogate-Based Measures
Prior works have introduced two surrogate-based measures:
ICE-Eval: ICE introduces its own evaluation measure (termed ICE-Eval). ICE-Eval’s surrogate assumes the U-CBEM’s projection operation $P$ has a reconstruction function $\tilde{\mathcal{P}}^{-1}:\mathbb{R}^{K}\to\mathcal{H}$. Their surrogate then is:
$s_{i}(.)$ $=f_{i}(\tilde{\mathcal{P}}^{-1}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (4)
where $f_{i}(.)$ denotes the $i$’th output from function $f(.)$. ICE-Eval introduces two metrics. Both are only applicable to classification models.
$ICE_{1}=\frac{1}{|\mathcal{V}|}\sum_{\mathbf{x},t\in\mathcal{V}}\frac{|y_{t}-\hat{y}_{t}|}{|y_{t}|}$ (5)
where $t$ is the index corresponding to the groundtruth class of $\mathbf{x}$. The second metric is:
[…]
C-SHAP-Eval: C-SHAP introduces its own evaluation measure (termed C-SHAP-Eval). Its surrogate learns a two-layer perceptron $\psi_{\text{C-SHAP}}:\mathbb{R}^{K}\to\mathcal{H}$:
$s_{i}(.)$ $=f_{i}(\psi_{\text{C-SHAP}}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (7)
The metric they define is:
$CSHAP_{1}=\frac{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\hat{\mathbf{p}})\}-a_{r}}{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\mathbf{p})\}-a_{r}}$ (8)
where $a_{r}$ is the accuracy from random predictions.
ICE-Eval and C-SHAP-Eval fail all three of our desidarata. (1) A critical piece of both surrogates is to reconstruct the model’s original embedding, and this reconstruction may be non-linear. This step makes the explanation indirectly related to the model’s output, placing a larger burden on the human interpreter. Additionally, C-SHAP introduces learnable parameters via $\psi_{\text{C-SHAP}}$. (2) Neither surrogate depends on concept importances ($A_{i}^{\prime}s$). Thus, unfaithful $A_{i}$’s will not alter faithfulness scores. (3) Error is measured with respect to the predicted or groundtruth class, ignoring errors on the remainder of the output distribution.
5 Method: Surrogate Faithfulness (SURF)
Surrogate Faithfulness (SURF) has two components. A simple surrogate maps the explanation to the model’s output (Fig. 1, Step 3); then, outputs from the surrogate and the original model are compared (Fig. 1, Step 4).
5.1 The SURF Surrogate
The final linear layer $\mathbf{F}=\begin{bmatrix}\mathbf{F}_{1}&\cdots&\mathbf{F}_{C}\end{bmatrix}$ operates on representation $\mathbf{H}$ to give the output $\mathbf{y}\in\mathcal{Y}$. In terms of their components, $\mathbf{H}=\begin{bmatrix}\mathbf{h}_{1}&\cdots&\mathbf{h}_{HW}\end{bmatrix}^{T}$ and $\mathbf{F}_{i}=\begin{bmatrix}\mathbf{f}_{i,1}&\cdots&\mathbf{f}_{i,HW}\end{bmatrix}^{T}\in\mathbb{R}^{(HW)\times D}$. Thus, output $y_{i}$ of $\phi$ is given by:
$y_{i}$ $=\sum_{j=1}^{HW}\mathbf{h}_{j}^{T}\mathbf{f}_{i,j}=\sum_{j=1}^{HW}\mathbf{h_{j}}^{T}\frac{\mathbf{f}_{i,j}}{{\|\mathbf{f}_{i,j}\|}_{2}}{\|\mathbf{f}_{i,j}\|}_{2}$ (9)
$\triangleq\sum_{j=1}^{HW}\mathbf{h}_{j}^{T}\mathbf{v}_{i,j}\alpha_{i,j}\quad\text{where }{\|\mathbf{v}_{i,j}\|}_{2}=1\quad\forall\ i,j$
where $\alpha_{i,j}={\|\mathbf{f}_{i,j}\|}_{2}$. In words, the model’s linear layer projects each embedding $\mathbf{h}_{j}$ onto a learned, class-specific direction (or CAV) $\mathbf{f}_{i,j}$. The projection is scaled by the norm (or importance) $\alpha_{i,j}$ of the learned direction. If the model uses global pooling to reduce the final embedding of $\mathbf{x}$ to a single vector, the summation over $j$ above can be omitted.
We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we input the concept basis representation and $A_{i}$’s from any U-CBEM. The surrogate is:
$\hat{y_{i}}$ $=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
where $\mathcal{P}(.)_{k}$ denotes the $k$’th element. Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer, allowing its linear surrogate to fully represent the model’s computation.
Table 3: Measure-over-measure comparison. We compare surrogate-based measures across three explanation settings (Perfect, Rand Imp, and Full Rand). We expect evaluations in the Perfect setting to give perfect faithfulness scores, and we expect progressively worse faithfulness evaluations as we increase randomness (just importances in Rand Imp and fully random in Full Rand). We include the SURF metrics (MAE and EMD) along with other metrics used by prior works. The SURF measure behaves as expected, whereas C-SHAP-Eval reports higher faithfulness in the Full Rand setting and ICE-Eval reports perfect faithfulness scores in the Rand Imp setting.
Surrogate Top-1 (%) Rank Corr Norm L1 MAE EMD Learnable Params.
Perfect C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
ICE-Eval 100 1.00 0.00 0.00 0.000 0
[…]
Rand Imp C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
ICE-Eval 100 1.00 0.00 0.00 0.000 0
[…]
Full Rand C-SHAP-Eval (CEL) 86.9 0.42 12.2 14.9 0.249 1M
C-SHAP-Eval (L1) 7.7 0.53 1.29 0.729 0.853 1M
ICE-Eval 3.4 0.07 1.00 1.66 0.862 0
[…]
Figure 1: Surrogate Faithfulness (SURF) overview. (1) Probe the model with a dataset of inputs to obtain intermediate embeddings from the penultimate layer $L$. (2) A training subset of the embeddings is used to fit the U-CBEM, discovering CAVs and concept importances for all classes $C$. (3) A test set of embeddings (along with the learned U-CBEM parameters) are passed to the surrogate to be mapped to predictions. (4) The SURF metrics compare surrogate and model outputs to measure faithfulness.
5.2 The SURF Metrics
The SURF metrics measure the difference in model and surrogate outputs (Fig. 1, Step 4). For classification models, this means the logits. However, a model’s logit-space is unconstrained, varying drastically between models. This makes logit-space metrics difficult to interpret. Furthermore, metrics in the logit-space dilute the importance of the predicted class by aggregating errors over all classes. On the other hand, the probability-space (after softmax) normalizes the logits, allowing for a bounded, interpretable metric. However, low error in the probability-space does not always imply low error in the logit-space. Due to the softmax, the probability-space emphasizes the predicted class and diminishes the other classes, so one could achieve a low error by accurately reproducing the predicted class and ignoring the others. Since each space addresses the other’s flaws, SURF measures errors in both spaces.
Let $\phi$ be a classification model, where $\mathbf{y}\in\mathbb{R}^{C}$ denotes class logits. To measure errors in the logit-space, we use the mean absolute error (MAE) between the logits:
$MAE=\frac{1}{|\mathcal{V}|C}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|y_{i}-\hat{y}_{i}|$ (11)
where $\mathcal{V}$ is the set of inputs used to evaluate for faithfulness. To measure errors in the probability-space, we find the Earth Mover’s Distance (EMD) using a constant distance cost. Letting $\mathbf{p}\in\mathbb{R}^{C}$ denote class probability scores,
$EMD=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
For a U-CBEM to be faithful, its explanations must have low EMD and MAE errors. For regression, we only use MAE.
In contrast to prior surrogate-based approaches, Surrogate Faithfulness (SURF) meets all of our desidarata. (1) It introduces a simple surrogate with no additional learnable parameters: the output for any class $i$ is a linear combination of the concept basis, weighted by the concept importances ($A_{i}$). The surrogate is not reconstruction-based; it tries to predict the next linear representation (e.g., class logits), instead of reconstructing embedding $\mathbf{h}$. Thus, only concepts useful for the prediction are needed, which may only be a subset of the concepts needed for reconstructing $\mathbf{h}$ from $\mathcal{P}(\mathbf{h};V)$. (2) SURF’s surrogate incorporates $A_{i}$, allowing inaccuracies in $A_{i}$ to affect faithfulness scores. (3) SURF defines two metrics $d$ that measure errors across the entire output space, instead of specific classes, to comprehensively assess faithfulness. These differences are summarized in Tab. 2.
6 Experiments
We perform experiments to: (1) empirically prove that SURF is a better faithfulness measure than prior ones; (2) benchmark prior U-CBEMs using SURF across three tasks, finding that many state-of-the-art U-CBEMs have poor faithfulness; (3) analyze the impact of the number of concepts on faithfulness; (4) study the robustness of U-CBEMs in adversarial settings, a novel scenario that has been unexplored in prior evaluations.
Table 4: Benchmarking U-CBEM faithfulness. We apply SURF to evaluate explanations from prior U-CBEMs in three tasks. Along with the SURF metrics (MAE and EMD), we report other metrics to serve as a comparison. We find that most prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
(a) Object classification
(ResNet-50)
U-CBEM MAE ($\downarrow$) EMD ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
CDISCO 2.29 0.957 0.9 0.000
ICE 3.06 0.438 93.1 0.222
CRAFT 1.70 0.844 78.9 0.484
C-SHAP 2.40 0.587 41.7 0.461
MCD 1.42 0.079 97.0 0.808
(b) Multi-attribute prediction
MAE ($\downarrow$) Attr-Acc (%) ($\uparrow$)
6.77 50.7
5.55 76.1
6.87 19.6
7.75 51.0
2.83 96.6
(c) Age
MAE ($\downarrow$)
32.6
6.1 Measure-over-Measure Comparison
We validate SURF against ICE-Eval and C-SHAP-Eval. As there is no “groundtruth” of faithfulness to compare evaluation results to, we test the frameworks in three manufactured settings, where we know the relative faithfulness of the explanations a priori. Specifically, we look at faithfulness results for a perfect explanation and two randomized explanations. To generate the Perfect explanation, observe that the most faithful explanation of the model is the model itself. Thus, we set $\{A_{i},V_{i}\}_{i=1}^{C}$ according to the weights of the linear classification layer. In our second setting (termed Rand Imp), we keep the perfect CAVs fixed and randomly sample the concept importances. Our final setting (termed Full Rand) additionally randomly samples the CAVs. In the Perfect setting, we expect to obtain no faithfulness error, regardless of the evaluation measure. Then, we expect to obtain increasingly less faithful scores as we progressively increase randomness (from Rand Imp to Full Rand).
We evaluate the faithfulness in these three settings on an ImageNet-pretrained ResNet-50 [16] finetuned on the Caltech-101 dataset [21]. Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq. 6 and ICE’s normalized L1 logit error (Norm L1) defined in Eq. 5. As a complementary measure to EMD, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs. This measure ensures that fine-grained changes that result in a different prediction ordering are penalized but ignores differences in magnitudes. Finally, we measure surrogate complexity with the number of learnable parameters the evaluation measure introduces. Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics. More details in Appendix Sec. A.
Results are reported in Tab. 3, averaged over 10 seeds. In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected. However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations. In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because they do not use concept importances, so even random concept importances have the same result as the Perfect setting. Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics. Finally, in the Full Rand setting, C-SHAP-Eval unintuitively reports an improvement in faithfulness compared to the Perfect setting. ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected. Of the tested faithfulness measures, SURF is the only one that behaves as expected, as validated by all metrics.
Next, we turn our attention to the metrics. We highlight a failure case of Top-1 by looking at C-SHAP-Eval (CEL)’s results in the Full Rand setting. The Top-1 score is strong; however, the other metrics (especially EMD) reveal that the reconstructed output has severe errors on the remainder of the output distribution. Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1), despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits. We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across other classes (as measured by MAE) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
As shown in the three settings presented, SURF is able to appropriately measure the faithfulness of three toy U-CBEMs. This contrasts with other approaches and metrics, which fail to yield satisfactory measurements in the same three cases. The remainder of the paper uses SURF to evaluate prior U-CBEMs.
6.2 Benchmarking U-CBEMs
We evaluate five prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, and MCD). We benchmark U-CBEMs on a varied range of tasks with a diverse range of models to demonstrate the applicability of SURF. Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression. In all tasks, each U-CBEM discovers 5 CAVs (or subspaces for MCD) per output. CAVs and importances are found on the training set, and the resulting explanations are evaluated on the test set. Dataset, finetuning, and U-CBEM implementation details are in Sec. B. Good-faith efforts were made to adapt prior U-CBEMs (that were developed for object classification) to new tasks.
We evaluate on three tasks. (1) Object Classification: We finetune a ResNet-50 on Caltech-101 and report MAE, EMD, Top-1, and Rank Corr. (2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [31] to predict the presence/absence of attributes in the CelebA dataset [22]. We report MAE and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1. (3) Age Estimation: Given a human face, we finetune a ViT [6] to predict their age (in years) on the UTK-Face dataset [45]. Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used. C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs. For this task, we report MAE.
Benchmark results are reported in Tab. 4. Aside from MCD, we observe that no prior U-CBEM is faithful to the original model, on any task. In the Object Classification task, the most faithful U-CBEM (after MCD) achieves the lowest EMD of 0.438, denoting large errors in the probability-space which stem from errors in the logit-space. Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task. CDISCO exhibits an average error of 32.6 years in the Age Estimation task.
We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding. Crucially, the class-specific CAVs are found on images from only that class, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will result in large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors. This issue is exacerbated for CRAFT & ICE, which non-linearly project to the concept space. (2) Except for ICE and MCD, the concept importance scores do not satisfy a logit-based completeness criterion (note, C-SHAP satisfies an accuracy-based completeness criterion). Thus, the concept importances do not faithfully reflect the impact of the concept representation on the model’s outputs. Though MCD also finds class-specific concepts, MCD achieves superior faithfulness by allowing a concept to be the span of many CAVs (a linear subspace); thus, it is able to faithfully explain a larger part of the model’s reasoning with the same number of concepts.
Figure 2: We fit U-CBEMs on Sec. 6.2’s Object Classification task with an increasing number of concepts and evaluate them with SURF. Some U-CBEMs do not improve as the number of concepts increase. U-CBEMs that improve quickly plateau, indicating an “elbow” phenomenon.
6.3 Faithfulness vs. Parsimony
The most important hyperparameter for any U-CBEM is the number of concepts to be discovered. Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation. However, presenting users with explanations containing too many concepts may be uninterpretable. The cognitive psychology literature has shown that humans can hold a limited number of items in their working memory at a time [5, 25]. Thus, simple, or parsimonious, explanations are preferred, which can communicate the bulk of the model’s reasoning in a few concepts. Instead of setting this hyperparameter arbitrarily, something commonly done in prior U-CBEMs (10 in [44], 25 in [9, 13]), we use SURF to intelligently analyze this tradeoff.
Using the Object Classification task from Sec. 6.2, we fit U-CBEMs with an increasing number of concepts. Each U-CBEM is evaluated using the SURF measure, visualized in Fig. 2. We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts. Only ICE and MCD uniformly improve as the number of concepts increases, and they exhibit an “elbow” phenomenon, marking a natural choice for the number of concepts.
Table 5: Novel setting: faithfulness on adversarial images. U-CBEMs are fixed from Sec. 6.2’s Object Classification task. The original model is adversarially attacked, and we report the mean paired difference in MAE and EMD (adversary - original) on 100 randomly sampled images. Any statistically insignificant difference is omitted (with “-”). U-CBEMs are generally not robust to adversarial attack; all of them exhibit an increase in MAE.
U-CBEM MAE ($\downarrow$) EMD ($\downarrow$)
CDISCO ++0.634 ++0.039
ICE ++0.730 -
CRAFT ++0.411 ++0.073
C-SHAP ++0.631 -0.195
MCD ++0.339 -0.038
6.4 Faithfulness in Adversarial Settings
Prior work has restricted their faithfulness evaluations to explanations on images from the same dataset the U-CBEM was trained on, as we did in Sec. 6.2. Such evaluations do not test U-CBEMs in challenging situations; yet, we believe these situations are the ones where having faithful explanations are the most critical. One such scenario is adversarial attack, which we test here. We take on the role of an adversary who has white-box access to the original model, but is unaware of the post-hoc explanation. If the U-CBEM is robust to adversarial attack, then we expect the U-CBEM to faithfully explain the model when it is under adversarial attack, allowing stakeholders to be aware of the adversarial attack. We design a Projected Gradient Descent (PGD) attack [24] to maximize the probability distribution difference of the original model on the input image and the adversary (details in Appendix Sec. C). The PGD attack is repeated on a 100 randomly sampled images from the training set for U-CBEMs fit in the Object Classification task from Sec. 6.2.
Tab. 5 reports the mean paired difference in MAE and EMD (adversarial - original). We use a one-sided, paired t-test on per-image differences to determine statistically significant increases in error (if $p<0.01$). We find that the MAE error increase is significant for all U-CBEMs, while EMD only significantly increases for CDISCO and CRAFT. We conclude that all U-CBEMs are less robust in the logit space, with ICE being the most vulnerable. The logit-space errors only propagate to the probability space for certain U-CBEMs (CDISCO & CRAFT). In certain cases (e.g., C-SHAP), the probability-space error may actually decrease on adversarial examples. We anticipate that U-CBEMs may be more vulnerable to adversarial attacks specifically designed to fool the explainer (instead of the model); as demonstrated under a simple PGD attack, SURF makes it possible to study such scenarios.
7 Conclusion
This paper organizes prior faithfulness measures, discusses their limitations, and proposes SURF to accurately measure faithfulness. We validate SURF compared to other measures in a measure-over-measure study, and we use SURF to find that several existing U-CBEMs produce unfaithful explanations. We apply SURF to downstream settings (number of concepts selection and adversarial robustness) to demonstrate its utility. We urge future works to adopt SURF to comprehensively evaluate faithfulness and compare to prior works.
Supplementary Material
This supplementary section provides information not included in the main text due to space constraints. It includes:
  1. 1.
    Sec. A: Implementation details for the surrogate justification experiment (Sec. 6.1).
  2. 2.
    Sec. B: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
  3. 3.
    Sec. C: Experimental details for the adversarial attack experiment done in Sec. 6.4.
A Implementation Details for Measure-over-Measure Comparison
This section contains details for the measure-over-measure comparison conducted in Sec. 6.1.
A.1 Experimental Setting Details
In the Perfect setting, we set CAVs and importances according to the weights of the final classification layer. Specifically, each class has 1 CAV ($K=1$). The CAV $\mathbf{v_{c,1}}$ for class $c$ is set to $\frac{\mathbf{f}_{c}}{{\|\mathbf{f}_{c}\|}_{2}}$, where $\mathbf{f}_{c}$ is class $c$’s classification vector. The CAV’s importance $\alpha_{c,1}$ is set to ${\|\mathbf{f}_{c}\|}_{2}$.
For the random settings, we sample each dimension of the CAVs independently from a standard normal distribution, and then make the CAV unit norm. We sample concept importances uniformly from $[0,1)$.
A.2 Surrogate Implementation
Recall that U-CBEMs define their own mechanism $\mathcal{P}:\mathcal{H}\to\mathbb{R}^{K}$ to project embeddings to the concept basis. For these experiments, since we are not using any U-CBEM and instead directly setting the CAVs and importances, we define $\mathcal{P}(\mathbf{h}_{j},V_{i})_{k}=\mathbf{h}_{j}^{T}\mathbf{v}_{i,k}$, where $i$ denotes the class, $k$ denotes the CAV, and $j$ selects the spatial embedding. Recall in this case that $K=1$ and there is only one embedding (obtained from global pooling).
A.2.1 SF Surrogate (Ours)
Following Sec. 5.1 and Eq. 10, we set the parameters of the surrogate with the concepts and importances obtained from the setting. This surrogate has no learnable parameters and is exactly equal to the original linear. Note that the embeddings are linearly projected onto the CAVs to find the concept representation.
A.2.2 ICE Surrogate
ICE’s surrogate takes the concept representation and reconstructs the embedding, as specified by the U-CBEM. It then passes through the final classification layer to obtain outputs. Let $\mathbf{p_{i}}\in\mathbb{R}^{K}$ denote the concept representation for class $i$ obtained from embedding $\mathbf{h}$. The surrogate is:
$y_{i}$ $=\mathbf{f}_{i}^{T}\psi_{ICE}(\mathbf{p_{i}})+b_{i}$ (13)
where $\psi_{ICE}(.)$ denotes the reconstruction, as specified by the U-CBEM. Since we are not making use of any U-CBEM in the measure-over-measure comparison, we use linear operations to reconstruct the embedding from the concept representation. That is, $\psi(\mathbf{p_{i}})=\sum_{k=1}^{K}p_{i,k}\mathbf{v}_{i,k}$.
A.2.3 C-SHAP Surrogate
C-SHAP’s surrogate defines a two-layer MLP $\psi_{\text{C-SHAP}}(.)$ to reconstruct the embedding, and then passes through the final classification layer to obtain outputs. The surrogate is:
$y_{i}$ $=\mathbf{f}_{i}^{T}\psi_{\text{C-SHAP}}(\mathbf{p_{i}})+b_{i}$ (14)
Since C-SHAP’s surrogate contains learnable parameters. Following [39], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities. We train the surrogate using the Adam optimizer [19] and a starting learning rate of $0.1$. The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus. We stop training when the learning rate is reduced below $10^{-7}$.
B Additional Details for Benchmarking U-CBEM Faithfulness
Sec. 6.2 evaluates prior U-CBEMs, finding that they produce unfaithful explanations. This section gives experimental and implementation details for this evaluation.
B.1 Model Details
Object Classification: Our evaluation uses a ResNet-50. We use the ImageNet pre-trained weights and initialize the final classification layer with random weights. The model is finetuned (conv4 and all subsequent layers) on the Caltech-101 training set, achieving a test accuracy of approximately 92.75%. The model is trained without data augmentation and using cross entropy loss.
Multi-Attribute Prediction: We finetune an ImageNet pre-trained MobileNetV2 on the CelebA dataset for the task of binary attribute prediction. We initialize the final prediction linear layer with random weights, and finetune all layers of the model. We achieve an attribute prediction accuracy of 91.55%. The model is trained with data augmentation and binary cross-entropy loss.
Age Regression: We finetune an ImageNet pre-trained Vision Transformer (ViT) on the UTK-Face dataset for the task of age regression. We choose the ViT-B/16 variant for our experiment. We initialize the final prediction linear layer with random weights, and finetune all weights above (not including) the 9th layer. The model achieves a mean average error of 5.14 on the test set. The model is trained with data augmentation and using L1 Loss.
B.2 Dataset Details
We learn each class’s concepts and importances using images of the class from the training set. Importantly, the label of each image is determined using the original model’s prediction (instead of the ground truth label). We do this because we are interested in explaining why the model predicts a specific label, regardless of whether the prediction is correct or not. After learning the U-CBEMs, we evaluate them on a test set.
Object Classification: We split Caltech-101’s dataset into train and test sets using 80/20 splits. While splitting the data, we maintain the same class distribution present in the original dataset in both the training and testing splits. To handle Caltech-101’s class imbalance, we assemble a balanced test set; this ensures that we evaluate for faithfulness equally across all classes.
Multi-Attribute Prediction: We use CelebA’s pre-defined train and test sets. We do not make any modifications to the test set. Since the training set is too large for many U-CBEMs, we take a random subset of training data to use for finding U-CBEM parameters. The random subset is equivalent in size to the test set.
Age Regression: We split the dataset into train and test sets using 80/20 splits. While splitting the data, we discretize the age distribution into bins of 10-years, and maintain this distribution across both the training and testing splits. As UTK-Face does not have any classes, all U-CBEMs find CAVs on all embeddings.
B.3 U-CBEM Details
This section specifies implementation details for each U-CBEM that was benchmarked.
B.4 CDISCO
The open-source implementation provided by the authors was used to evaluate this approach. CDISCO uses SVD to discover CAVs. Then, a gradient-based approach is used to calculate the importance of each CAV. The code used to compute the gradients was not provided in the repository, so we re-implemented it. Embeddings are linearly projected to the concept space.
Object Classification: Due to a lack of clarity, we had to make one implementation choice. Specifically, when finding the importance (using a gradient calculation), it was unclear if we should use images from all classes or only images of the CAV’s class. We opted for the latter, based on the code and the example provided by the authors.
Multi-Attribute Prediction: CDISCO was extended to this task by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently
Age Regression: The open-source implementation only considered the case of multi-class classification. According to the paper and the code used for multi-class classification, we re-implemented CDISCO to work in the case of single-output regression.
B.5 ICE
The open-source implementation provided by the authors was used to evaluate this approach. ICE uses NMF to find CAVs. Due to the potentially large number of embeddings, we use the batched-version of NMF. Concept importance is found using the TCAV method proposed in [17]. Embeddings are projected to the concept space using sci-kit learn’s transform method. ICE was extended to multi-attribute prediction by reformulating the task as independent binary classification tasks. Thus, CAVs and concept importances were learned for each task independently.
B.6 CRAFT
The open-source implementation provided in the Xplique toolbox [8] (by the same authors) was used to evaluate this approach. CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance. Embeddings are projected to the concept space using the toolbox’s built-in transform method.
B.7 ConceptSHAP
ConceptSHAP differs from previous U-CBEMs because it discovers CAVs globally (i.e., class-agnostic). Thus, we need a way to associate discovered CAVs to the concepts. To do so, we discover a (larger than $K$) set of CAVs and assign them a per-class concept importance score (through Shapley Values). Then, the top $K$ CAVs by importance magnitude are selected to be the CAVs for a given class. For all experiments, we initially discover 100 CAVs. This number is arbitrary, but note that the Shapley Value computation scales exponentially with the number of initially discovered CAVs, which motivated our choice. With more CAVs, we expect ConceptSHAP to obtain higher faithfulness at the expense of a higher computational cost. Embeddings are linearly projected to the concept space. For age regression, we only discover $K$ concepts (since there are no classes). For multi-attribute prediction, we repeat the ConceptSHAP procedure for each attribute, considering each attribute prediction as a binary classification task; we only discover $K$ concepts, since each task is a binary classification.
The authors provide an open-source implementation in TensorFlow, so we instead use a third-party re-implementation in PyTorch. Several modifications are made to this implementation:
  1. 1.
    Added in a learnable two-layer MLP to map back to the representation space (as done in the original paper)
  2. 2.
    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([23]) and is class-specific.
  3. 3.
    Fixed the implementation of the completeness score
B.8 MCD
Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces. In [35], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace. To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$. Then, any embedding $\mathbf{h}$ can be uniquely decomposed onto the concept basis as:
$\mathbf{h}=\sum_{l=1}^{n_{c}+1}\sum_{j=1}^{d^{\,l}}p_{j}^{i,l}\,\mathbf{v}_{j}^{i,l}$ (15)
Thus, we stack all $p_{j}^{i,l}$ into the concept representation $\mathbf{p_{i}}$ for class $i$. Since $C^{i,\perp}$ is not interpretable, we set elements of $\mathbf{p_{i}}$ corresponding to $C^{i,\perp}$ to 0; thus, the reconstruction of $\mathbf{h}$ is done using interpretable concepts presented to the user during an explanation. We set concept importance $\alpha_{j}^{i,l}=\mathbf{f}_{i}^{T}\mathbf{v}_{j}^{i,l}$, which directly relates to the local and global concept relevance defined in MCD and allows for faithful reconstruction of the output using our surrogate.
Since MCD is computationally expensive when trying to discover concepts on a large number of embeddings, we randomly sample 10,000 embeddings if the number of embeddings extracted for a given class is above 10,000. We use the open-source implementation provided by the authors.
C Adversarial Attack Details
We conducted a white-box adversarial attack using Projected Gradient Descent (PGD) [24] to evaluate the robustness of U-CBEMs under adversarial perturbations. The attack aims to maximize the EMD between the probability outputs of the model on the original image and the adversary.
A PGD attack with 100 iterations, step size $\alpha=1$, and perturbation bound $\epsilon=0.1$ was implemented to iteratively perturb input images while maximizing the EMD loss. The final perturbed image with the highest EMD loss (across all iterations) was used for evaluation. The original implementation of PGD assumes that the images are normalized from [0,1] to define the perturbation bound. Our images were standardized (following typical ImageNet processing), so we empirically selected an $\epsilon$ that resulted in minor perceptible changes to the reference image.

### v3

@@
 Abstract
-Post-hoc, unsupervised concept-based explanation methods (U-CBEMs) translate a vision model’s internal reasoning into human-understandable concepts, leading to interpretable explanations.
-However, we find that many state-of-the-art (SOTA) U-CBEMs are not faithful: their concepts seem interpretable but fail to reproduce the model’s predictions.
-We argue that this deficiency has gone unnoticed due to fragmented evaluation—each paper proposes its own faithfulness measure, with no measure-over-measure comparison or broad benchmarking.
-We close this gap by (i) organizing prior metrics in a unified framework, discussing their limitations, and identifying desiderata for a faithfulness measure; (ii) introducing the Surrogate Faithfulness (SURF) measure, which quantifies faithfulness via the predictive loss of a surrogate that maps explanations to the model’s outputs; and (iii) delivering the first comprehensive U-CBEM faithfulness benchmark across diverse tasks and architectures.
-In a controlled setting, SURF outperforms prior faithfulness measures in measure-over-measure comparisons, and applying SURF to SOTA U-CBEMs reveals that many visually appealing U-CBEMs are surprisingly unfaithful.
-We demonstrate SURF applicability in two downstream settings—(i) faithfulness versus the number of concepts used in the explanation and (ii) U-CBEM robustness to adversarial attacks—underscoring SURF’s value as a reliable faithfulness measure.
+Deep vision models perform input-output computations that are hard to interpret.
+Concept-based explanation methods (CBEMs) increase interpretability by re-expressing parts of the model with human-understandable semantic units, or concepts.
+Checking if the derived explanations are faithful—that is, they represent the model’s internal computation—requires a surrogate that combines concepts to compute the output.
+Simplifications made for interpretability inevitably reduce faithfulness, resulting in a tradeoff between the two.
+State-of-the-art unsupervised CBEMs (U-CBEMs) have reported increasingly interpretable concepts, while also being more faithful to the model.
+However, we observe that the reported improvement in faithfulness artificially results from either (1) using overly complex surrogates, which introduces an unmeasured cost to the explanation’s interpretability, or (2) relying on deletion-based approaches that, as we demonstrate, do not properly measure faithfulness.
+We propose Surrogate Faithfulness (SURF), which (1) replaces prior complex surrogates with a simple, linear surrogate that measures faithfulness without changing the explanation’s interpretability and (2) introduces well-motivated metrics that assess loss across all output classes, not just the predicted class.
+We validate SURF with a measure-over-measure study by proposing a simple sanity check—explanations with random concepts should be less faithful—which prior surrogates fail.
+SURF enables the first reliable faithfulness benchmark of U-CBEMs, revealing that many visually compelling U-CBEMs are not faithful.
 Code to be released.
 1 Introduction
-Deep learning models have delivered state-of-the-art results across diverse tasks, yet their internal reasoning remain difficult to interpret [12], which is especially problematic in high-stakes domains such as healthcare and finance.
+Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [Gerlings2020ReviewingTN], which is especially problematic in high-stakes domains such as healthcare and finance.
 This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior.
-Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that accurately reflect the model’s internal reasoning and allow humans to reliably predict the model’s output.
+Two properties characterize explanation quality: interpretability—explanations that humans can understand—and faithfulness—explanations that reflect the model’s internal computation.
 XAI methods must balance the natural tension between interpretability and faithfulness.
-While interpretability is typically assessed with human studies, faithfulness can be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model.
-Because explanations are inevitably lossy, some discrepancy between the surrogate and the model is expected; a faithfulness measure must therefore specify both (i) an appropriate surrogate and (ii) a metric for the induced loss.
-One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model.
-Since the explanation and the model are merged, these approaches eliminate the need for faithfulness checks; however, they often underperform black-box models by reducing model complexity (capacity) for explainability.
-In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution—commonly found through gradients or perturbations [38, 30, 23, 28, 7].
-For images, these attributions are visualized as pixel-level heatmaps.
-Although such explanations are intuitive, attributions are usually unfaithful, since they are estimated from noisy gradients in deep neural networks or from unrealistic perturbations on the input space [1, 11, 32, 15].
-The heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [26, 4].
-Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [10, 3, 17, 46].
-To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss aspects of the model’s reasoning, raising faithfulness concerns [29].
-To avoid annotation, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output.
-We argue that, in practice, U-CBEMs have prioritized interpretability at the expense of faithfulness, yielding explanations that look compelling but are surprisingly unfaithful.
-Despite the importance of faithful explanations, we find that all prior works introduce new faithfulness measures, with no measure-over-measure comparison and sparse cross-U-CBEM evaluations.
-Thus, there is a lack of consensus on faithfulness measures for U-CBEMs, limiting our understanding of how faithful current approaches are.
-This work serves to address this gap. Our contributions are to:
+While interpretability is assessed with human studies, faithfulness must be evaluated by defining a surrogate that maps the explanation to the model’s outputs and measuring the loss between the surrogate and model.
+Since explanations are inevitably lossy, some discrepancy is expected.
+For certain XAI methods, defining a surrogate is challenging, so alternative measures, or proxies, attempt to evaluate faithfulness using different criteria.
+One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [protopnet, senn].
+Merging the explanation and the model eliminates the need for faithfulness checks; however, these methods often underperform black-box models by reducing model complexity (capacity) for explainability.
+In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [wang2024gradient, lime, shapley, Petsiuk2018rise, sobol].
+For images, attributions are visualized as pixel-level heatmaps.
+Although such explanations are intuitive, developing faithfulness measures has been challenging, since many attribution methods do not specify a surrogate to reconstruct the model’s output from the explanation.
+Thus, proxies—such as checking if important pixels come from foreground regions or looking for significant model degradation after deleting important pixels—are used to study faithfulness.
+Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [Adebayo2018SanityCF, Frye2020ShapleyManifold, Sixt2019WhenEL, Hase2021TheOP].
+The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [Nguyen2021TheEO, colin2022what].
+Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [net2vec, netdiss, kimTCAV, zhou2014object].
+To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [Ramaswamy2022OverlookedFI].
+To avoid these issues, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output.
+Since CBEMs operate on the model’s intermediate representations, defining a surrogate is more natural, paving the way for proper faithfulness evaluations.
+However, a closer look shows that U-CBEMs have adopted inadequate faithfulness measures.
+We discover that existing evaluations rely on complex surrogates, which allow U-CBEMs to simultaneously show users interpretable explanations and report high faithfulness, yet the explanations do not clearly lead to the model’s output.
+Furthermore, U-CBEMs adapt popular deletion-based proxies from the attribution literature, yet Sec.
+4.1 outlines serious unresolved limitations that prevent them from properly measuring faithfulness.
+Finally, each work that proposes a new U-CBEM also evaluates it with a new faithfulness measure, with no measure-over-measure comparison, indicating a lack of consensus on faithfulness measures in the field.
+We argue that these issues have misled us into believing that current U-CBEMs are faithful. To shed light on this, we make the following contributions:
   1. 1.
-    Organize prior U-CBEM faithfulness measures under a common framework, allowing us to discuss limitations in current measures and identify appropriate desidarata for a faithfulness measure.
+    Organize prior U-CBEM faithfulness measures under a common framework, allowing us to discuss their limitations and identify appropriate desiderata.
   2. 2.
-    Propose Surrogate Faithfulness (SURF), a faithfulness measure satisfying the desidarata.
-We conduct a measure-over-measure comparison to empirically validate our faithfulness measure against prior measures.
+    Propose Surrogate Faithfulness (SURF), a faithfulness measure satisfying the desiderata.
+Our measure-over-measure comparison checks if faithfulness decreases as the explanation is increasingly randomized; only SURF passes this check.
   3. 3.
     Conduct the first, comprehensive faithfulness benchmark of current U-CBEMs across a variety of tasks and model architectures.
-Most SOTA U-CBEMs that were thought to be faithful are found not to be.
+SURF shows that SOTA U-CBEMs previously evaluated to be faithful are not.
   4. 4.
-    Extend SURF to two downstream applications.
-First, studying faithfulness as a function of the number of concepts shows that some U-CBEMs do not get more faithful as they discover more concepts.
-Second, analyzing the faithfulness of U-CBEMs when the model is under adversarial attack reveals robustness issues.
+    Leverage SURF to provide a selection criterion for the number of concepts U-CBEMs discover, improving on prior work that sets this hyperparameter arbitrarily.
 2 Preliminaries
-We first introduce notation that will be used in the remainder of the paper, as well as preliminaries to familiarize the reader with the field.
+We first introduce notation used throughout the paper, as well as preliminaries to familiarize the reader with the field.
 Background: Model $\phi:\mathcal{X}\to\mathcal{Y}$ maps from an input space $\mathcal{X}\subseteq\mathbb{R}^{P}$ to an output space $\mathcal{Y}\subseteq\mathbb{R}^{C}$.
@@
 Let $g:\mathcal{X}\to\mathcal{H}$ and $f:\mathcal{H}\to\mathcal{Y}$.
-Thus, $\phi(\mathbf{X})=f(g(\mathbf{X}))$.
-Let $g(\mathbf{x})=\mathbf{h}\in\mathcal{H}$ represent the embedding of $\mathbf{x}$ (e.g., image patch), and $g(\mathbf{X})=\mathbf{H}\in\mathbb{R}^{(HW)\times D}$ denotes the vectorized application of $g$ on $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{HW}\end{bmatrix}^{T}$, where $H,W$ are the spatial dimensions.
-Let $E$ be an explanation function, and let $\mathbf{e}=E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image).
+Thus, $\mathbf{y}=\phi(\mathbf{X})=f(g(\mathbf{X}))$.
+Let $g(\mathbf{x})=\mathbf{h}\in\mathcal{H}$ represent the embedding of $\mathbf{x}$ (e.g., image patch), and $g(\mathbf{X})=\mathbf{H}\in\mathbb{R}^{(HW)\times D}$ denote the vectorized application of $g$ on $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\cdots&\mathbf{x}_{HW}\end{bmatrix}^{T}$, where $H,W$ are the spatial dimensions.
+Let $E$ be an explanation function, and let $E(\mathbf{X})\in\mathcal{E}$ represent the explanation for an $\mathbf{X}\in\mathcal{X}$ (e.g., image).
 Let $s:\mathcal{E}\to\mathcal{Y}$ be a surrogate.
 Motivated by Ribeiro et al.
-[30], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$.
+[lime], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$.
 Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
 $Faith(E;d,s)=\int_{\mathcal{X}}d(\phi(\mathbf{X}),s(E(\mathbf{X})))d\mathbf{X}$ (1)
-U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and their concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$.
-Note that while $K$ can vary per output, all prior works find the same number of CAVs for each output.
+U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$.
+Note that while $K$ can vary per output, all prior works find the same $K$ per output.
 U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$.
@@
 Following prior works, we study faithfulness in the final layer, where the spatial dimensions $H=W=1$, so $\mathbf{H}=\mathbf{h}$.
-Given that U-CBEMs operate on $g(X)$, the faithfulness measure can be simplified to:
+Given that U-CBEMs operate on $g(\mathbf{X})$, the faithfulness measure can be simplified to:
 $Faith_{\text{U-CBEM}}(\mathcal{P},\{V_{i}\}_{i=1}^{C},\{A_{i}\}_{i=1}^{C};d,s)=$ (3)
 $\int_{\mathcal{H}}d(f(\mathbf{h}),\{s(\mathcal{P}(\mathbf{h};V_{i}),A_{i})\}_{i=1}^{C})d\mathbf{h}$
+Figure 1: Framework.
+U-CBEM faithfulness measures compare (using metric $d$) the output $\mathbf{y}$ of the original model $\phi(.)=f(g(.))$ with the output $\mathbf{\hat{y}}$ from the explanation.
+To obtain the explanation, U-CBEMs transform intermediate representation $\mathbf{h}\in\mathcal{R}^{D}$ to the concept representation through concept projection $\mathcal{P}(\mathbf{h})$ and provide an accompanying concept importance AA.
+The explanation is then passed through surrogate $s$ to obtain $\mathbf{\hat{y}}$.
+Deletion-based proxies (right) observe model degradation after performing deletion in a deletion space.
+Surrogate-based measures (left) do not manipulate the computation; instead, they introduce a surrogate to directly approximate Eq.
+3.
 3 Related Works
-3.1 Unsupervised Concept-Based Explanations
+Unsupervised Concept-Based Explanation Methods.
 Wang et al.
-[37] initiated this line of work, finding CAVs by K-Means clustering intermediate embeddings computed on a dataset.
-Each cluster was associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches.
+[Wang2015UnsupervisedLO] find CAVs by K-Means clustering intermediate embeddings computed on a dataset.
+Each cluster is associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches.
 Ghorbani et al.
-[13] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate.
+[Ghorbani2019ACE] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate.
 To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al.
-[39] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [41] to find concept importance.
-ICE [44] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation.
+[ConceptSHAP] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [Young1985MonotonicSO] to find concept importance.
+ICE [Zhang2020InvertibleCE] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation.
 The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA).
 Building on this, Fel et al.
-[9] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance.
-CDISCO [14] uses the singular value decomposition (SVD) to find CAVs and a gradient-based method for finding concept importance.
-Recently, MCD [35] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
-Table 1: Deletion-Based Measures: We organize deletion-based measures introduced by prior works according to their differing factors.
-The factors we use are Perturbation Space, Deletion Strategy, and how the results are presented (Y-axis and X-axis).
-Notice how all prior works introduce a different faithfulness measure, with no measure-over-measure comparison.
-Method Perturbation Space Deletion Strategy Y-axis X-axis
-ACE Pixel Constant (grey) Accuracy # Deleted Concepts
-CDISCO (M1) Pixel Constant (grey) # Classes with $80\%$ accuracy loss # Deleted Concepts
-MCD Pixel Inpainting Accuracy Fraction of Inpainted Pixels
-CRAFT Concept Constant (zero) Change in Class Logit Score # Deleted Concepts
-CDISCO (M2) Weight Constant (zero) Accuracy # Deleted Concepts
-3.2 Concept Bottleneck Models
-U-CBEMs are not to be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model.
-These methods [20, 18, 42, 27] train a model to first predict a set of interpretable concepts (e.g., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification).
-By doing so, CBMs fully decomposes the model’s prediction into interpretable concepts, and exposing the concept representation allows users to intervene (e.g., increase the strength of the “red” concept to help the model recognize the “stop sign”).
+[Fel2022CRAFTCR] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance.
+CDISCO [graziani2023concept] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance.
+MCD [vielhaben2023multidimensional] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
+Finding issues with MCD’s interpretability, HU-MCD [hu-mcd] incorporates Segment Anything Model [SAM] to discover more interpretable concepts.
+Table 1: Prior deletion-based proxies (top) & surrogate-based measures (bottom) are organized according to their differing factors: Deletion Space (DS), Deletion Method (DM), Metric, Surrogate, and Concept Projection (CP).
+Each prior work uses a different proxy or measure, with no measure-over-measure comparison.
+Method DS DM Metric Surrogate CP
+ACE Pixel Constant (grey) Class Accuracy Original Model Identity
+MCD Pixel Inpainting Class Accuracy Original Model Identity
+HU-MCD Pixel Masking Class Accuracy Original Model Identity
+CDISCO (M1) Pixel Constant (grey) # Classes with $80\%$ Accuracy Loss Original Model Identity
+CDISCO (M2) Weight Constant (zero) Class Accuracy Original Model Identity
+CRAFT Concept Constant (zero) Class Logit Reconstruct $\to$ Original Model U-CBEM
+ICE Normalized Target Class L1 Logit Loss U-CBEM Reconstruct $\to$ Original Model U-CBEM
+Top-1 Accuracy
+C-SHAP Normalized Top-1 Accuracy MLP Reconstruct $\to$ Original Model U-CBEM
+Concept Bottleneck Models.
+CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model.
+These methods [pmlr-cbm-koh20a, prob-cbm, posthoc-cbm, labelfree-cbm] train a model to first predict a set of interpretable concepts (i.e., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification).
+By doing so, CBMs fully decompose the prediction into interpretable concepts.
 While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts.
 As CBMs are inherently explainable, the notion of faithfulness does not extend to them.
-This work will only discuss faithfulness as it pertains to U-CBEMs.
+This work discusses faithfulness only for U-CBEMs.
 4 Unifying U-CBEM Faithfulness Measures
-We provide a comprehensive review of all faithfulness measures that were introduced by the works mentioned in Sec.
-3.1.
-We find that each work uses a different measure for faithfulness, indicating a lack of unification in the field and preventing fair comparisons of faithfulness across different U-CBEMs.
-As part of our unified framework, we place prior faithfulness measures into two categories: 1) deletion-based and 2) surrogate-based.
-We detail the dimensions across which prior deletion-based measures differ; this is followed by a discussion on their limitations, which raise questions on the validity of past assessments.
-Then, we similarly characterize prior surrogate-based measures and demonstrate their drawbacks.
-4.1 Deletion-Based Measures
-Deletion-based measures evaluate faithfulness by observing the effect of deleting a concept on the model’s output.
-Intuitively, these measures assert that removing the most important concepts will result in the greatest model degradation.
-After obtaining CAVs and their importances (for a class), concepts are removed (Deletion Strategy) from the input’s Perturbation Space in decreasing order of importance, and the effect on the model is measured after each deletion.
-Results are visualized by plotting the effect on the model (Y-axis) with respect to some measure of how many concepts were deleted (X-axis).
-Prior works make different choices along each italicized dimension mentioned.
-Perturbation Space: defines the space in which concepts are removed.
-This is either the image space, the model’s weight space, or the concept space.
-If making removals in the pixel or the weight space, the U-CBEM must have some mechanism to associate the concept to this space (e.g., which pixels in the image represent the concept).
-Deletion Strategy: defines how concepts are removed. Concepts are usually removed from the Perturbation Space by setting the concept to a baseline value (e.g., 0).
-Y-axis: defines what is measured after each deletion. Typically, this is accuracy (e.g., in object classification).
-X-axis: represents how much “information” was deleted. This is typically the number of deleted concepts.
-To understand the differences between prior deletion measures, Tab.
-1 organizes prior measures according to these dimensions.
-ACE [13] deletes concepts in the image space by setting associated pixels to a constant value (Constant Deletion).
-They measure the model’s prediction accuracy after each deletion across a dataset.
-CRAFT [9] transforms the model’s embeddings into a concept representation.
-It makes Constant Deletions in the concept space by setting the concept to zero; then, they reconstruct the embedding from the perturbed concept representation.
-They measure the change in the corresponding class’s logit score.
-CDISCO [14] has two measures.
-The first measure (M1) makes Constant Deletions in the pixel space and measures the number of classes that have a greater than 80% accuracy degradation.
-The second measure (M2) deletes concepts by zeroing out parameters in the model’s weight space that are associated with the concept, and they plot classification accuracy.
-MCD [35] deletes concepts in the pixel space by using inpainting.
-They measure model classification accuracy with respect to the fraction of inpainted pixels, unlike all prior measures that use the number of deleted concepts.
-4.1.1 Limitations of Deletion-Based Measures
-There are two notable issues with deletion-based metrics that prevent them from accurately evaluating faithfulness.
+Each U-CBEM mentioned in Sec.
+3 introduces a different faithfulness measure, indicating a lack of consensus and preventing fair comparisons of faithfulness across U-CBEMs.
+We unify faithfulness measures under a common framework, shown in Fig.
+1.
+Any evaluation measure compares the output of the original model with the output from the explanation (Metric).
+To obtain the output from the explanation, a Concept Projection transforms the intermediate representation to a concept representation, which is then passed through a Surrogate.
+Deletion-based proxies assert that removing the most important concepts will result in the greatest model degradation; more degradation is interpreted as a signal for higher faithfulness.
+Alternatively, surrogate-based measures avoid changing the input and introduce a surrogate to directly approximate faithfulness as in Eq.
+3.
+4.1 Deletion-Based Proxies
+To perform deletion, deletion-based proxies must remove concepts (Deletion Method) from the input’s Deletion Space in decreasing order of importance.
+Deletion Space: Where concepts are removed. This is either the image space, the model’s weight space, or the concept space (shown in Fig. 1 (right)).
+Deletion Method: How concepts are removed from the Deletion Space, usually by setting the concept to a baseline value (e.g., 0).
+Except for CRAFT, proxies set the concept projection to identity and the surrogate to $f(.)$ from the original model.
+Prior proxies (organized in Tab.
+1) make different choices along each italicized factor mentioned above.
+ACE deletes concepts in the image space by setting associated pixels to a constant value (Constant Deletion).
+They measure the model’s classification accuracy after each deletion across a dataset.
+MCD measures classification accuracy after deleting concepts in the pixel space with inpainting.
+HU-MCD similarly measures classification accuracy but performs deletions through a masking strategy which simulates running the model on the irregularly-shaped input, better ignoring masked regions.
+CDISCO has two measures.
+The first measure (M1) makes Constant Deletions on pixels and measures the number of classes whose accuracy degrades by more than 80%.
+The second measure (M2) deletes concepts by zeroing out model parameters associated with the concept, and they report classification accuracy.
+CRAFT uses the U-CBEMs defined $\mathcal{P}$ to transform $\mathbf{h}$ to a concept representation.
+It makes Constant Deletions in the concept space by setting the concept to zero; then, it reconstructs $\mathbf{h}$ from the perturbed concept representation and measures the change in the corresponding class’s logit score.
+There are two notable issues with deletion-based proxies that prevent them from accurately evaluating faithfulness.
 First, it is unclear how to delete a concept.
 Concepts are usually deleted by Constant Deletion, commonly with a baseline of 0.
-The baseline must entirely delete the presence of the concept without affecting other concepts.
-However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [11].
+This baseline must entirely delete the presence of the concept without affecting other concepts.
+However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [Frye2020ShapleyManifold].
 Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
-Second, there are no guarantees that representations in the Perturbation Space will stay on manifold after deletion.
-The feature attribution literature [11, 2, 34, 33, 40] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold.
+Second, there are no guarantees that representations in the Deletion Space will stay on manifold after deletion.
+The feature attribution literature [Frye2020ShapleyManifold, akata2022manifold, taufiq2023manifold, Slack2019FoolingLA, yeh2022threading] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold.
 In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features.
-Thus, measuring feature importance with off-manifold inputs can lead to incorrect explanations [11].
-More seriously, these explanations can be adversarially attacked; in [33], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs.
-Such issues can be mitigated by generating explanations using on-manifold input perturbations but requires accurate modeling of the input distribution [11, 34, 36, 43, 15].
-Deletions in the concept or weight space manipulate intermediate representations, so the same concerns hold.
-Thus, we argue that deletion metrics have significant unresolved issues that introduce unreliability in their faithfulness evaluations, so we avoid using these approaches.
+Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [Frye2020ShapleyManifold].
+More seriously, these explanations can be adversarially attacked; in [Slack2019FoolingLA], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs.
+Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [Frye2020ShapleyManifold, taufiq2023manifold, vu2022emap, zaher2024manifold, Hase2021TheOP].
+These problems are not unique to feature attribution; they extend to deletions in the concept and weight space, which make similar manipulations in the model’s intermediate representation space.
+Thus, we argue that deletion-based proxies are unreliable, and we avoid using or benchmarking against them.
 4.2 Surrogate-Based Measures
-Surrogate-based measures evaluate faithfulness by approximating (Eq 3).
-Any surrogate-based measure must make two choices: 1) what the surrogate $s$ is, and 2) what the metric $d$ is.
-We will introduce prior faithfulness measures, and then we discuss their limitations.
-4.2.1 Desidarata
+Any surrogate-based measure must choose the surrogate $s$ and the metric $d$.
+For these measures, the input image, and thus the intermediate representation, is the same in both paths of Fig.
+1 (left), and the concept projection is defined by the U-CBEM being evaluated.
+4.2.1 Desiderata
 The faithfulness definition in Eq.
-3 does not constrain the surrogate $s$, but there are certain properties our surrogate should have.
+3 does not constrain the surrogate $s$, but there are certain desiderata our surrogate should satisfy.
 First, $s$ should be as simple as possible.
-Recognize that $s$ reflects the mental computation any human interpreter may be expected to do when trying to map the explanation onto the model’s prediction.
-An interpretable explanation that requires a complex $s$ is effectively useless; it will not give the human interpreter any meaningful insight into how the model reaches its prediction.
-Second, the $s$ should incorporate all components of the explanation.
-For example, U-CBEMs explain with both $V_{i}$ and $A_{i}$, so both should be used by the surrogate.
+Recognize that $s$ reflects the mental computation any human interpreter is expected to do when trying to connect the explanation to the model’s prediction.
+An interpretable explanation that requires a complex $s$ just shifts the interpretation issue downstream; the human interpreter still lacks insight into how the model reaches its prediction.
+Second, $s$ should incorporate all components of the explanation.
+For example, U-CBEMs explain with both $V_{i}$ and $A_{i}$, so both should be used by $s$.
 If a component is not included, its impact on faithfulness cannot be measured.
-Third, we desire metric(s) $d$ that rewards explanations that can closely reconstruct all model outputs (e.g.
-across all classes), instead of only the model’s predicted class.
-This directly follows from Eq.
-3.
-Table 2: Surrogate-Based Measures: We compare prior surrogate-based measures introduced in ICE and C-SHAP to the proposed Surrogate Faithfulness (SURF) measure.
-According to our desidarata, we require surrogates that are simple (e.g., not reconstruction-based and not requiring additional parameters), incorporate concept importances into the surrogate, and measure errors across all outputs.
-Prior works fail to meet these desidarata (except for ICE not requiring additional parameters).
-We mark in green and red the properties that meet our desidarata.
+Third, we desire metric(s) $d$ that rewards explanations that closely reconstruct all of $\mathbf{y}$ (e.g.
+across all classes), not just specific (e.g., predicted) classes.
+4.2.2 Prior Surrogate-Based Measures
+Two prior works have introduced surrogate-based measures:
+ICE-Eval: ICE introduces its own evaluation measure (we term as ICE-Eval).
+ICE-Eval’s surrogate assumes the U-CBEM’s projection operation $P$ has a reconstruction function $\tilde{\mathcal{P}}^{-1}:\mathbb{R}^{K}\to\mathcal{H}$.
+Their surrogate then is:
+$\mathbf{\hat{y}}_{i}=s_{i}(.)$ $=f_{i}(\tilde{\mathcal{P}}^{-1}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (4)
+where $f_{i}(.)$ denotes the $i$’th output from function $f(.)$. ICE-Eval uses two metrics (only applicable to classification):
+$ICE_{1}=\frac{1}{|\mathcal{V}|}\sum_{\mathbf{x},t\in\mathcal{V}}\frac{|y_{t}-\hat{y}_{t}|}{|y_{t}|}$ (5)
+where $t$ is the index corresponding to $\mathbf{x}$’s groundtruth class and $\mathcal{V}$ is the set of inputs used to evaluate faithfulness.
+Letting $\mathbf{p}$ denote class probability scores, the second metric is:
+[…]
+C-SHAP-Eval: C-SHAP introduces its own evaluation measure (we term as C-SHAP-Eval). Its surrogate learns a two-layer perceptron $\psi_{\text{C-SHAP}}:\mathbb{R}^{K}\to\mathcal{H}$:
+$\mathbf{\hat{y}}_{i}=s_{i}(.)$ $=f_{i}(\psi_{\text{C-SHAP}}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (7)
+The metric they define is:
+$CSHAP_{1}=\frac{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\hat{\mathbf{p}})\}-a_{r}}{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\mathbf{p})\}-a_{r}}$ (8)
+where $a_{r}$ is the accuracy from random predictions.
+ICE-Eval and C-SHAP-Eval fail all three of our desiderata.
+(1) A critical piece of both surrogates is to reconstruct the model’s original embedding, and this reconstruction may be non-linear.
+Additionally, C-SHAP introduces non-linearity and learnable parameters via $\psi_{\text{C-SHAP}}$.
+The overly complex surrogates place a large burden on the human interpreter, which is not reflected in prior interpretability and faithfulness evaluations.
+(2) Neither surrogate depends on concept importances ($A_{i}^{\prime}s$).
+Thus, unfaithful $A_{i}$’s will not alter faithfulness scores.
+(3) Error is measured only on the predicted or groundtruth class, ignoring errors on the remainder of the output distribution.
+Table 2: Surrogate-based desiderata We compare prior surrogate-based measures introduced in ICE and C-SHAP to the proposed Surrogate Faithfulness (SURF) measure.
+According to our desiderata, we require surrogates that are simple (e.g., not reconstruction-based and not requiring additional parameters), incorporate concept importances into the surrogate, and measure errors across all outputs.
+Prior measures largely fail to meet these desiderata.
+We mark in green and red the properties that meet our desiderata.
 A ✓  or ✗  denotes that the method does or does not have the property.
@@
 Errors on All Outputs?
-4.2.2 Prior Surrogate-Based Measures
-Prior works have introduced two surrogate-based measures:
-ICE-Eval: ICE introduces its own evaluation measure (termed ICE-Eval).
-ICE-Eval’s surrogate assumes the U-CBEM’s projection operation $P$ has a reconstruction function $\tilde{\mathcal{P}}^{-1}:\mathbb{R}^{K}\to\mathcal{H}$.
-Their surrogate then is:
-$s_{i}(.)$ $=f_{i}(\tilde{\mathcal{P}}^{-1}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (4)
-where $f_{i}(.)$ denotes the $i$’th output from function $f(.)$. ICE-Eval introduces two metrics. Both are only applicable to classification models.
-$ICE_{1}=\frac{1}{|\mathcal{V}|}\sum_{\mathbf{x},t\in\mathcal{V}}\frac{|y_{t}-\hat{y}_{t}|}{|y_{t}|}$ (5)
-where $t$ is the index corresponding to the groundtruth class of $\mathbf{x}$. The second metric is:
+Table 3: Measure-over-measure comparison.
+We compare surrogate-based measures across three explanation settings (Perfect, Rand Imp, and Full Rand).
+We expect evaluations in the Perfect setting to give perfect faithfulness scores, and we expect progressively worse faithfulness evaluations as we increase randomness (just importances in Rand Imp and fully random in Full Rand).
+We include $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ along with metrics used by prior works.
+The SURF metrics behaves as expected, whereas C-SHAP-Eval reports higher faithfulness in the Full Rand setting and ICE-Eval reports perfect faithfulness scores in the Rand Imp setting.
+Surrogate Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$) Norm L1 ($\downarrow$) $\text{{SURF}}_{\text{MAE}}$ ($\downarrow$) $\text{{SURF}}_{\text{EMD}}$ ($\downarrow$) Params Learnt ($\downarrow$) FLOPs ($\downarrow$)
+Perfect C-SHAP-Eval (CEL) 9.02 -0.02 1.27 1.97 0.865 1M 205M
+C-SHAP-Eval (L1) 6.13 0.08 2.15 0.54 0.883 1M 205M
+ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
 […]
-C-SHAP-Eval: C-SHAP introduces its own evaluation measure (termed C-SHAP-Eval). Its surrogate learns a two-layer perceptron $\psi_{\text{C-SHAP}}:\mathbb{R}^{K}\to\mathcal{H}$:
-$s_{i}(.)$ $=f_{i}(\psi_{\text{C-SHAP}}(\mathcal{P}(\mathbf{h};V_{i})))\;\;\forall\;i$ (7)
-The metric they define is:
-$CSHAP_{1}=\frac{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\hat{\mathbf{p}})\}-a_{r}}{\sum_{\mathbf{x},t\in\mathcal{V}}\mathds{1}\{t=argmax(\mathbf{p})\}-a_{r}}$ (8)
-where $a_{r}$ is the accuracy from random predictions.
-ICE-Eval and C-SHAP-Eval fail all three of our desidarata.
-(1) A critical piece of both surrogates is to reconstruct the model’s original embedding, and this reconstruction may be non-linear.
-This step makes the explanation indirectly related to the model’s output, placing a larger burden on the human interpreter.
-Additionally, C-SHAP introduces learnable parameters via $\psi_{\text{C-SHAP}}$.
-(2) Neither surrogate depends on concept importances ($A_{i}^{\prime}s$).
-Thus, unfaithful $A_{i}$’s will not alter faithfulness scores.
-(3) Error is measured with respect to the predicted or groundtruth class, ignoring errors on the remainder of the output distribution.
+Rand Imp C-SHAP-Eval (CEL) 9.02 -0.02 1.27 1.97 0.865 1M 205M
+C-SHAP-Eval (L1) 6.13 0.08 2.15 0.54 0.883 1M 205M
+ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
+[…]
+Full Rand C-SHAP-Eval (CEL) 97.6 0.02 181.7 168.2 0.125 1M 205M
+C-SHAP-Eval (L1) 6.1 0.08 3.59 1.721 0.883 1M 205M
+ICE-Eval 3.3 0.00 1.00 3.17 0.882 0 614K
+[…]
 5 Method: Surrogate Faithfulness (SURF)
-Surrogate Faithfulness (SURF) has two components.
-A simple surrogate maps the explanation to the model’s output (Fig.
-1, Step 3); then, outputs from the surrogate and the original model are compared (Fig.
-1, Step 4).
+Surrogate Faithfulness (SURF) introduces a simple, linear surrogate and two metrics; taken together, they allow SURF to satisfy the desiderata.
 5.1 The SURF Surrogate
@@
 In terms of their components, $\mathbf{H}=\begin{bmatrix}\mathbf{h}_{1}&\cdots&\mathbf{h}_{HW}\end{bmatrix}^{T}$ and $\mathbf{F}_{i}=\begin{bmatrix}\mathbf{f}_{i,1}&\cdots&\mathbf{f}_{i,HW}\end{bmatrix}^{T}\in\mathbb{R}^{(HW)\times D}$.
-Thus, output $y_{i}$ of $\phi$ is given by:
+Thus, output $y_{i}$ of $\phi$ is given by (up to the bias term):
 $y_{i}$ $=\sum_{j=1}^{HW}\mathbf{h}_{j}^{T}\mathbf{f}_{i,j}=\sum_{j=1}^{HW}\mathbf{h_{j}}^{T}\frac{\mathbf{f}_{i,j}}{{\|\mathbf{f}_{i,j}\|}_{2}}{\|\mathbf{f}_{i,j}\|}_{2}$ (9)
@@
 If the model uses global pooling to reduce the final embedding of $\mathbf{x}$ to a single vector, the summation over $j$ above can be omitted.
-We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we input the concept basis representation and $A_{i}$’s from any U-CBEM. The surrogate is:
-$\hat{y_{i}}$ $=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
+We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we use the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
+$\mathbf{\hat{y}}_{i}$ $=s_{i}(.)=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
 where $\mathcal{P}(.)_{k}$ denotes the $k$’th element.
-Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer, allowing its linear surrogate to fully represent the model’s computation.
-Table 3: Measure-over-measure comparison.
-We compare surrogate-based measures across three explanation settings (Perfect, Rand Imp, and Full Rand).
-We expect evaluations in the Perfect setting to give perfect faithfulness scores, and we expect progressively worse faithfulness evaluations as we increase randomness (just importances in Rand Imp and fully random in Full Rand).
-We include the SURF metrics (MAE and EMD) along with other metrics used by prior works.
-The SURF measure behaves as expected, whereas C-SHAP-Eval reports higher faithfulness in the Full Rand setting and ICE-Eval reports perfect faithfulness scores in the Rand Imp setting.
-Surrogate Top-1 (%) Rank Corr Norm L1 MAE EMD Learnable Params.
-Perfect C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
-C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
-ICE-Eval 100 1.00 0.00 0.00 0.000 0
-[…]
-Rand Imp C-SHAP-Eval (CEL) 9.06 -0.18 0.94 1.72 0.842 1M
-C-SHAP-Eval (L1) 0.51 0.44 1.34 0.86 0.846 1M
-ICE-Eval 100 1.00 0.00 0.00 0.000 0
-[…]
-Full Rand C-SHAP-Eval (CEL) 86.9 0.42 12.2 14.9 0.249 1M
-C-SHAP-Eval (L1) 7.7 0.53 1.29 0.729 0.853 1M
-ICE-Eval 3.4 0.07 1.00 1.66 0.862 0
-[…]
-Figure 1: Surrogate Faithfulness (SURF) overview.
-(1) Probe the model with a dataset of inputs to obtain intermediate embeddings from the penultimate layer $L$.
-(2) A training subset of the embeddings is used to fit the U-CBEM, discovering CAVs and concept importances for all classes $C$.
-(3) A test set of embeddings (along with the learned U-CBEM parameters) are passed to the surrogate to be mapped to predictions.
-(4) The SURF metrics compare surrogate and model outputs to measure faithfulness.
+Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer.
+Thus, its surrogate fully represents the model’s computation, while greatly reducing complexity compared to prior faithfulness surrogates.
 5.2 The SURF Metrics
-The SURF metrics measure the difference in model and surrogate outputs (Fig.
-1, Step 4).
+The SURF metrics measure the difference in model and surrogate outputs.
 For classification models, this means the logits.
@@
 Since each space addresses the other’s flaws, SURF measures errors in both spaces.
-Let $\phi$ be a classification model, where $\mathbf{y}\in\mathbb{R}^{C}$ denotes class logits. To measure errors in the logit-space, we use the mean absolute error (MAE) between the logits:
-$MAE=\frac{1}{|\mathcal{V}|C}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|y_{i}-\hat{y}_{i}|$ (11)
-where $\mathcal{V}$ is the set of inputs used to evaluate for faithfulness.
-To measure errors in the probability-space, we find the Earth Mover’s Distance (EMD) using a constant distance cost.
-Letting $\mathbf{p}\in\mathbb{R}^{C}$ denote class probability scores,
-$EMD=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
-For a U-CBEM to be faithful, its explanations must have low EMD and MAE errors. For regression, we only use MAE.
-In contrast to prior surrogate-based approaches, Surrogate Faithfulness (SURF) meets all of our desidarata.
-(1) It introduces a simple surrogate with no additional learnable parameters: the output for any class $i$ is a linear combination of the concept basis, weighted by the concept importances ($A_{i}$).
-The surrogate is not reconstruction-based; it tries to predict the next linear representation (e.g., class logits), instead of reconstructing embedding $\mathbf{h}$.
-Thus, only concepts useful for the prediction are needed, which may only be a subset of the concepts needed for reconstructing $\mathbf{h}$ from $\mathcal{P}(\mathbf{h};V)$.
-(2) SURF’s surrogate incorporates $A_{i}$, allowing inaccuracies in $A_{i}$ to affect faithfulness scores.
+If $\phi$ is a classification model, then $\mathbf{y}$ denotes class logits. To measure errors in the logit-space, we use the mean absolute error between the logits:
+$\text{SURF}_{\text{MAE}}=\frac{1}{|\mathcal{V}|C}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|y_{i}-\hat{y}_{i}|$ (11)
+To measure errors in the probability-space, we find the Earth Mover’s Distance using a constant distance cost.
+$\text{SURF}_{\text{EMD}}=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
+A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors. For regression, we only use $\text{SURF}_{\text{MAE}}$.
+In contrast to prior surrogate-based measures, SURF meets all of our desiderata.
+(1) It introduces a simple surrogate with no additional learnable parameters: the output for any class $i$ is a linear combination of the concept representation, weighted by the concept importances ($A_{i}$).
+The surrogate is not reconstruction-based; it tries to linearly predict the next representation (e.g., class logits), instead of reconstructing embedding $\mathbf{h}$.
+Thus, only concepts useful for the prediction are needed, which may be a subset of the concepts needed for reconstructing $\mathbf{h}$ from $\mathcal{P}(\mathbf{h};V)$.
+(2) SURF’s surrogate incorporates $A_{i}$, so inaccuracies in $A_{i}$ impact faithfulness scores.
 (3) SURF defines two metrics $d$ that measure errors across the entire output space, instead of specific classes, to comprehensively assess faithfulness.
@@
 2.
-6 Experiments
-We perform experiments to: (1) empirically prove that SURF is a better faithfulness measure than prior ones; (2) benchmark prior U-CBEMs using SURF across three tasks, finding that many state-of-the-art U-CBEMs have poor faithfulness; (3) analyze the impact of the number of concepts on faithfulness; (4) study the robustness of U-CBEMs in adversarial settings, a novel scenario that has been unexplored in prior evaluations.
 Table 4: Benchmarking U-CBEM faithfulness.
 We apply SURF to evaluate explanations from prior U-CBEMs in three tasks.
-Along with the SURF metrics (MAE and EMD), we report other metrics to serve as a comparison.
-We find that most prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
+Along with the $\text{SURF}_{\text{MAE}}$ and $\text{SURF}_{\text{EMD}}$, we report other metrics to serve as a comparison.
+We find that prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
 (a) Object classification
 (ResNet-50)
-U-CBEM MAE ($\downarrow$) EMD ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
-CDISCO 2.29 0.957 0.9 0.000
-ICE 3.06 0.438 93.1 0.222
-CRAFT 1.70 0.844 78.9 0.484
-C-SHAP 2.40 0.587 41.7 0.461
-MCD 1.42 0.079 97.0 0.808
+U-CBEM $\text{{SURF}}_{\text{MAE}}$ ($\downarrow$) $\text{{SURF}}_{\text{EMD}}$ ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
+CDISCO 3.40 0.932 0.2 0.002
+ICE 3.33 0.628 98.9 0.093
+CRAFT 3.19 0.878 90.6 0.068
+C-SHAP 3.28 0.882 6.3 0.005
+MCD 2.60 0.426 99.4 0.145
+HU-MCD 1.97 0.384 99.7 0.149
 (b) Multi-attribute prediction
-MAE ($\downarrow$) Attr-Acc (%) ($\uparrow$)
+$\text{{SURF}}_{\text{MAE}}$ Attr-Acc (%) ($\uparrow$)
 6.77 50.7
@@
 (c) Age
-MAE ($\downarrow$)
+$\text{{SURF}}_{\text{MAE}}$
 32.6
+6 Experiments
+We first perform a measure-over-measure study with a basic sanity check: do faithfulness scores decrease as we progressively randomize the explanation?
+Only SURF passes this check.
+Then, we benchmark prior U-CBEMs across three tasks, revealing that SOTA methods are not faithful.
+Finally, we leverage SURF to analyze the trade-off between the number of concepts in an explanation and faithfulness.
 6.1 Measure-over-Measure Comparison
-We validate SURF against ICE-Eval and C-SHAP-Eval.
-As there is no “groundtruth” of faithfulness to compare evaluation results to, we test the frameworks in three manufactured settings, where we know the relative faithfulness of the explanations a priori.
-Specifically, we look at faithfulness results for a perfect explanation and two randomized explanations.
+The fundamental challenge in evaluating faithfulness measures is that there is no “groundtruth” of faithfulness to compare evaluation results to.
+To address this, we propose a simple sanity check, which introduces three manufactured settings where the relative faithfulness of explanations is known a priori.
+Then, we check if the faithfulness measure preserves the relative ordering of scores across settings.
+Concretely, we look at faithfulness results for a perfect explanation and two randomized explanations.
 To generate the Perfect explanation, observe that the most faithful explanation of the model is the model itself.
@@
 Then, we expect to obtain increasingly less faithful scores as we progressively increase randomness (from Rand Imp to Full Rand).
-We evaluate the faithfulness in these three settings on an ImageNet-pretrained ResNet-50 [16] finetuned on the Caltech-101 dataset [21].
+We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [resnet] finetuned on the Caltech-101 dataset [caltech-101].
 Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq.
@@
 5.
-As a complementary measure to EMD, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs.
-This measure ensures that fine-grained changes that result in a different prediction ordering are penalized but ignores differences in magnitudes.
-Finally, we measure surrogate complexity with the number of learnable parameters the evaluation measure introduces.
+As a complementary measure to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs.
+Rank Corr penalizes fine-grained changes that result in a different prediction ordering but ignores differences in magnitudes.
+Finally, we measure surrogate complexity by its learnable parameters and FLOPs needed to obtain the output from the concept representation.
 Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics.
-More details in Appendix Sec.
+More details are in Appendix Sec.
 A.
@@
 However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations.
-In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because they do not use concept importances, so even random concept importances have the same result as the Perfect setting.
+In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because their surrogates do not use concept importances, so using random concept importances has the same evaluation result as the Perfect setting.
 Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics.
@@
 ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected.
-Of the tested faithfulness measures, SURF is the only one that behaves as expected, as validated by all metrics.
 Next, we turn our attention to the metrics.
-We highlight a failure case of Top-1 by looking at C-SHAP-Eval (CEL)’s results in the Full Rand setting.
-The Top-1 score is strong; however, the other metrics (especially EMD) reveal that the reconstructed output has severe errors on the remainder of the output distribution.
-Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1), despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits.
-We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across other classes (as measured by MAE) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
-As shown in the three settings presented, SURF is able to appropriately measure the faithfulness of three toy U-CBEMs.
-This contrasts with other approaches and metrics, which fail to yield satisfactory measurements in the same three cases.
-The remainder of the paper uses SURF to evaluate prior U-CBEMs.
+We highlight a failure case of Top-1 by looking at SURF’s results in the Rand Imp setting.
+The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the reconstructed output has severe errors on the remainder of the output distribution.
+Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits.
+We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across all classes (as measured by $\text{SURF}_{\text{MAE}}$) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
+Of the faithfulness measures, only SURF passes the sanity check while having the lowest surrogate complexity (fewest FLOPs and no learnable parameters).
+The remainder of the paper uses SURF to study prior U-CBEMs.
 6.2 Benchmarking U-CBEMs
-We evaluate five prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, and MCD).
-We benchmark U-CBEMs on a varied range of tasks with a diverse range of models to demonstrate the applicability of SURF.
+We evaluate six prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD).
+We benchmark U-CBEMs on a varied range of tasks and models to demonstrate the applicability of SURF.
 Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression.
-In all tasks, each U-CBEM discovers 5 CAVs (or subspaces for MCD) per output.
-CAVs and importances are found on the training set, and the resulting explanations are evaluated on the test set.
-Dataset, finetuning, and U-CBEM implementation details are in Sec.
+In all tasks, U-CBEMs discover 5 CAVs (or subspaces for MCD and HU-MCD) per output.
+CAVs and importances are found on the training set; the resulting explanations are evaluated on the test set.
+Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Sec.
 B.
-Good-faith efforts were made to adapt prior U-CBEMs (that were developed for object classification) to new tasks.
 We evaluate on three tasks.
-(1) Object Classification: We finetune a ResNet-50 on Caltech-101 and report MAE, EMD, Top-1, and Rank Corr.
-(2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [31] to predict the presence/absence of attributes in the CelebA dataset [22].
-We report MAE and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1.
-(3) Age Estimation: Given a human face, we finetune a ViT [6] to predict their age (in years) on the UTK-Face dataset [45].
+(1) Object Classification: We use the Caltech-101-finetuned ResNet-50 and report $\text{SURF}_{\text{MAE}}$, $\text{SURF}_{\text{EMD}}$, Top-1, and Rank Corr.
+HU-MCD requires architectural modifications, so we can only evaluate it on this task, which uses a compatible ResNet-50.
+(2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [Sandler2018MobileNetV2IR] to predict the presence/absence of attributes on the CelebA dataset [celeba].
+We report $\text{SURF}_{\text{MAE}}$ and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1.
+(3) Age Estimation: Given a human face, we finetune a ViT [VIT] to predict their age (in years) on the UTK-Face dataset [utkface].
 Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used.
 C-SHAP is omitted as it is only applicable to classification, and MCD is incompatible with ViTs.
-For this task, we report MAE.
 Benchmark results are reported in Tab.
 4.
-Aside from MCD, we observe that no prior U-CBEM is faithful to the original model, on any task.
-In the Object Classification task, the most faithful U-CBEM (after MCD) achieves the lowest EMD of 0.438, denoting large errors in the probability-space which stem from errors in the logit-space.
+In the evaluated settings, we observe that no prior U-CBEM is faithful to the original model.
+In the Object Classification task, the most faithful U-CBEM is HU-MCD with an $\text{SURF}_{\text{EMD}}$ of 0.384, denoting large errors in the probability-space.
 Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task.
@@
 We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding.
-Crucially, the class-specific CAVs are found on images from only that class, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will result in large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors.
-This issue is exacerbated for CRAFT & ICE, which non-linearly project to the concept space.
-(2) Except for ICE and MCD, the concept importance scores do not satisfy a logit-based completeness criterion (note, C-SHAP satisfies an accuracy-based completeness criterion).
-Thus, the concept importances do not faithfully reflect the impact of the concept representation on the model’s outputs.
-Though MCD also finds class-specific concepts, MCD achieves superior faithfulness by allowing a concept to be the span of many CAVs (a linear subspace); thus, it is able to faithfully explain a larger part of the model’s reasoning with the same number of concepts.
+Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors.
+This issue is exacerbated for CRAFT and ICE, which non-linearly project to the concept space.
+Though MCD and HU-MCD also find class-specific concepts, they achieve superior faithfulness by using concept subspaces; thus they faithfully explain more of the model with the same number of concepts.
+(2) Most methods do not satisfy a logit-based completeness criterion, preventing a simple reproduction of model outputs.
+Note that C-SHAP satisfies an accuracy-based completeness criterion and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be presented to the user in the explanation.
 Figure 2: We fit U-CBEMs on Sec.
-6.2’s Object Classification task with an increasing number of concepts and evaluate them with SURF.
+6.2’s Object Classification task with an increasing number of concepts and evaluate their faithfulness with SURF.
 Some U-CBEMs do not improve as the number of concepts increase.
-U-CBEMs that improve quickly plateau, indicating an “elbow” phenomenon.
+U-CBEMs that improve quickly plateau.
 6.3 Faithfulness vs. Parsimony
@@
 Intuitively, having more concepts should result in a more complete (i.e., faithful) explanation.
-However, presenting users with explanations containing too many concepts may be uninterpretable.
-The cognitive psychology literature has shown that humans can hold a limited number of items in their working memory at a time [5, 25].
-Thus, simple, or parsimonious, explanations are preferred, which can communicate the bulk of the model’s reasoning in a few concepts.
-Instead of setting this hyperparameter arbitrarily, something commonly done in prior U-CBEMs (10 in [44], 25 in [9, 13]), we use SURF to intelligently analyze this tradeoff.
+However, explanations containing too many concepts may be difficult to interpret.
+Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [2010TheMM, Miller1956TheMN].
+Thus, simple, or parsimonious, explanations are preferred, which communicate the bulk of the model’s computation with a few concepts.
+However, inadequate faithfulness measures make the faithfulness-parsimony tradeoff hard to analyze, so prior U-CBEMs commonly set this hyperparameter arbitrarily (10 in ICE, 25 in CRAFT & ACE).
+SURF allows us to intelligently analyze this tradeoff.
 Using the Object Classification task from Sec.
@@
 We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts.
-Only ICE and MCD uniformly improve as the number of concepts increases, and they exhibit an “elbow” phenomenon, marking a natural choice for the number of concepts.
-Table 5: Novel setting: faithfulness on adversarial images.
-U-CBEMs are fixed from Sec.
-6.2’s Object Classification task.
-The original model is adversarially attacked, and we report the mean paired difference in MAE and EMD (adversary - original) on 100 randomly sampled images.
-Any statistically insignificant difference is omitted (with “-”).
-U-CBEMs are generally not robust to adversarial attack; all of them exhibit an increase in MAE.
-U-CBEM MAE ($\downarrow$) EMD ($\downarrow$)
-CDISCO ++0.634 ++0.039
-ICE ++0.730 -
-CRAFT ++0.411 ++0.073
-C-SHAP ++0.631 -0.195
-MCD ++0.339 -0.038
-6.4 Faithfulness in Adversarial Settings
-Prior work has restricted their faithfulness evaluations to explanations on images from the same dataset the U-CBEM was trained on, as we did in Sec.
-6.2.
-Such evaluations do not test U-CBEMs in challenging situations; yet, we believe these situations are the ones where having faithful explanations are the most critical.
-One such scenario is adversarial attack, which we test here.
-We take on the role of an adversary who has white-box access to the original model, but is unaware of the post-hoc explanation.
-If the U-CBEM is robust to adversarial attack, then we expect the U-CBEM to faithfully explain the model when it is under adversarial attack, allowing stakeholders to be aware of the adversarial attack.
-We design a Projected Gradient Descent (PGD) attack [24] to maximize the probability distribution difference of the original model on the input image and the adversary (details in Appendix Sec.
-C).
-The PGD attack is repeated on a 100 randomly sampled images from the training set for U-CBEMs fit in the Object Classification task from Sec.
-6.2.
-Tab.
-5 reports the mean paired difference in MAE and EMD (adversarial - original).
-We use a one-sided, paired t-test on per-image differences to determine statistically significant increases in error (if $p<0.01$).
-We find that the MAE error increase is significant for all U-CBEMs, while EMD only significantly increases for CDISCO and CRAFT.
-We conclude that all U-CBEMs are less robust in the logit space, with ICE being the most vulnerable.
-The logit-space errors only propagate to the probability space for certain U-CBEMs (CDISCO & CRAFT).
-In certain cases (e.g., C-SHAP), the probability-space error may actually decrease on adversarial examples.
-We anticipate that U-CBEMs may be more vulnerable to adversarial attacks specifically designed to fool the explainer (instead of the model); as demonstrated under a simple PGD attack, SURF makes it possible to study such scenarios.
+ICE marginally improves as we add more concepts.
+Only MCD and HU-MCD uniformly improve as the number of concepts increases, and they exhibit a plateauing effect, marking a natural choice for the number of concepts.
+Interestingly, HU-MCD is more faithful than MCD only when both discover relatively few concepts; as the number of concepts grows, MCD becomes more faithful.
+This offers a more nuanced perspective on the findings from the HU-MCD paper, where they claimed superior faithfulness over MCD based on results from their deletion-based proxy.
 7 Conclusion
-This paper organizes prior faithfulness measures, discusses their limitations, and proposes SURF to accurately measure faithfulness.
-We validate SURF compared to other measures in a measure-over-measure study, and we use SURF to find that several existing U-CBEMs produce unfaithful explanations.
-We apply SURF to downstream settings (number of concepts selection and adversarial robustness) to demonstrate its utility.
-We urge future works to adopt SURF to comprehensively evaluate faithfulness and compare to prior works.
+This paper argues that we lack a clear view on how faithful current U-CBEMs really are, largely because of measures that deviate from the formal definition of faithfulness.
+In accordance with this definition, we develop desiderata for a faithfulness measure and propose SURF as a simple, principled measure for faithfulness.
+SURF passes a basic sanity check, reveals that state-of-the-art U-CBEMs are substantially unfaithful, and is a practical tool for studying the tradeoff between parsimony and faithfulness when choosing the number of concepts.
+We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores alongside their interpretability claims.
+[…]
 Supplementary Material
@@
   1. 1.
-    Sec. A: Implementation details for the surrogate justification experiment (Sec. 6.1).
+    Sec. A: Implementation details for the measure-over-measure study (Sec. 6.1).
   2. 2.
     Sec. B: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
-  3. 3.
-    Sec. C: Experimental details for the adversarial attack experiment done in Sec. 6.4.
 A Implementation Details for Measure-over-Measure Comparison
@@
 Recall that U-CBEMs define their own mechanism $\mathcal{P}:\mathcal{H}\to\mathbb{R}^{K}$ to project embeddings to the concept basis.
-For these experiments, since we are not using any U-CBEM and instead directly setting the CAVs and importances, we define $\mathcal{P}(\mathbf{h}_{j},V_{i})_{k}=\mathbf{h}_{j}^{T}\mathbf{v}_{i,k}$, where $i$ denotes the class, $k$ denotes the CAV, and $j$ selects the spatial embedding.
+For these experiments, since we are not using any U-CBEM and instead directly setting the CAVs and importances, we define $\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}=\mathbf{h}_{j}^{T}\mathbf{v}_{i,k}$, where $i$ denotes the class, $k$ denotes the CAV, and $j$ selects the spatial embedding.
 Recall in this case that $K=1$ and there is only one embedding (obtained from global pooling).
-A.2.1 SF Surrogate (Ours)
+A.2.1 SURF Surrogate (Ours)
 Following Sec.
 5.1 and Eq.
-10, we set the parameters of the surrogate with the concepts and importances obtained from the setting.
-This surrogate has no learnable parameters and is exactly equal to the original linear.
+10, we set the parameters of the surrogate with the obtained concepts and importances.
+This surrogate has no learnable parameters and is exactly equal to the original linear layer (up to the bias term).
 Note that the embeddings are linearly projected onto the CAVs to find the concept representation.
+After the surrogate reconstructs the output, we add the constant bias term before evaluation.
 A.2.2 ICE Surrogate
@@
 Since C-SHAP’s surrogate contains learnable parameters.
-Following [39], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities.
-We train the surrogate using the Adam optimizer [19] and a starting learning rate of $0.1$.
+Following [ConceptSHAP], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities.
+We train the surrogate using the Adam optimizer [kingma2014adam] and a starting learning rate of $0.1$.
 The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus.
 We stop training when the learning rate is reduced below $10^{-7}$.
+A.3 FLOPs Computation
+We treat multiply and add as 2 FLOPs. Let $K$ be the number of concepts, $C$ be the number of output classes, and $D$ by the dimensionality of the embedding space.
+SURF Surrogate: Given that the surrogate is a linear layer from concept representations to the output space (though it uses different concept representations for each class), the number of FLOPs required is $2KC$.
+ICE Surrogate: This surrogate reconstructs the original embedding (by a linear projection) from the concept representation, and then uses the original model (which is also a linear projection) to map to the output space.
+Thus, the number of FLOPs required is $CD(2K+1)$.
+C-SHAP Surrogate: This surrogate reconstructs the original embedding (with a two-layer MLP) from the concept representation, and then uses the original model (which is a linear projection) to map to the output space.
+Thus, the number of FLOPs required is $2C(H_{1}K+H_{1}D+D)$, where $H_{1}$ is the dimensionality of the MLP’s hidden layer.
 B Additional Details for Benchmarking U-CBEM Faithfulness
@@
 B.1 Model Details
-Object Classification: Our evaluation uses a ResNet-50.
+Object Classification: Our evaluation uses a ResNet-50, available from the timm library.
 We use the ImageNet pre-trained weights and initialize the final classification layer with random weights.
-The model is finetuned (conv4 and all subsequent layers) on the Caltech-101 training set, achieving a test accuracy of approximately 92.75%.
-The model is trained without data augmentation and using cross entropy loss.
+The model is finetuned (conv4 and all subsequent layers) on the Caltech-101 training set, achieving a test accuracy of approximately 95.61%.
+The model is trained without data augmentation and using cross entropy loss with label smoothing.
 Multi-Attribute Prediction: We finetune an ImageNet pre-trained MobileNetV2 on the CelebA dataset for the task of binary attribute prediction.
@@
 Due to the potentially large number of embeddings, we use the batched-version of NMF.
-Concept importance is found using the TCAV method proposed in [17].
+Concept importance is found using the TCAV method proposed in [kimTCAV].
 Embeddings are projected to the concept space using sci-kit learn’s transform method.
@@
 B.6 CRAFT
-The open-source implementation provided in the Xplique toolbox [8] (by the same authors) was used to evaluate this approach.
+The open-source implementation provided in the Xplique toolbox [fel2022xplique] (by the same authors) was used to evaluate this approach.
 CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance.
@@
   2. 2.
-    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([23]) and is class-specific.
+    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([shapley]) and is class-specific.
   3. 3.
@@
 Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces.
-In [35], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace.
+In [vielhaben2023multidimensional], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace.
 To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$.
@@
 We set concept importance $\alpha_{j}^{i,l}=\mathbf{f}_{i}^{T}\mathbf{v}_{j}^{i,l}$, which directly relates to the local and global concept relevance defined in MCD and allows for faithful reconstruction of the output using our surrogate.
-Since MCD is computationally expensive when trying to discover concepts on a large number of embeddings, we randomly sample 10,000 embeddings if the number of embeddings extracted for a given class is above 10,000.
+Since MCD is computationally expensive when trying to discover concepts on a large number of embeddings, we randomly sample 10,000 embeddings if there are greater than these many embeddings extracted for a given class.
 We use the open-source implementation provided by the authors.
-C Adversarial Attack Details
-We conducted a white-box adversarial attack using Projected Gradient Descent (PGD) [24] to evaluate the robustness of U-CBEMs under adversarial perturbations.
-The attack aims to maximize the EMD between the probability outputs of the model on the original image and the adversary.
-A PGD attack with 100 iterations, step size $\alpha=1$, and perturbation bound $\epsilon=0.1$ was implemented to iteratively perturb input images while maximizing the EMD loss.
-The final perturbed image with the highest EMD loss (across all iterations) was used for evaluation.
-The original implementation of PGD assumes that the images are normalized from [0,1] to define the perturbation bound.
-Our images were standardized (following typical ImageNet processing), so we empirically selected an $\epsilon$ that resulted in minor perceptible changes to the reference image.
+B.9 HU-MCD
+HU-MCD directly extends upon MCD, adopting the same framework of concept subspace, orthogonal complement, and unique decomposition onto the concept basis.
+The only material difference is in the concept discovery stage, where concepts are discovered on representations arising from image segments, instead of from the feature map.
+Therefore, we make the same adaptations as MCD for HU-MCD.
+We use the open-source implementation provided by the authors.

### v4

@@
 Simplifications made for interpretability inevitably reduce faithfulness, resulting in a tradeoff between the two.
-State-of-the-art unsupervised CBEMs (U-CBEMs) have reported increasingly interpretable concepts, while also being more faithful to the model.
+State-of-the-art unsupervised CBEMs (U-CBEMs) are seemingly more interpretable, while also being more faithful to the model.
 However, we observe that the reported improvement in faithfulness artificially results from either (1) using overly complex surrogates, which introduces an unmeasured cost to the explanation’s interpretability, or (2) relying on deletion-based approaches that, as we demonstrate, do not properly measure faithfulness.
@@
 SURF enables the first reliable faithfulness benchmark of U-CBEMs, revealing that many visually compelling U-CBEMs are not faithful.
-Code to be released.
+Code is released.
 1 Introduction
-Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [Gerlings2020ReviewingTN], which is especially problematic in high-stakes domains such as healthcare and finance.
+Deep learning models have delivered state-of-the-art (SOTA) results across diverse tasks, yet their internal computation remain difficult to interpret [17], which is especially problematic in high-stakes domains such as healthcare and finance.
 This gap has given rise to explainable AI (XAI), whose goal is to produce explanations of model behavior.
@@
 For certain XAI methods, defining a surrogate is challenging, so alternative measures, or proxies, attempt to evaluate faithfulness using different criteria.
-One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [protopnet, senn].
+One broad family of XAI methods constructs inherently interpretable models, injecting an explanatory structure directly into the model [6, 3].
 Merging the explanation and the model eliminates the need for faithfulness checks; however, these methods often underperform black-box models by reducing model complexity (capacity) for explainability.
-In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [wang2024gradient, lime, shapley, Petsiuk2018rise, sobol].
+In contrast, feature attribution methods express a trained model’s outputs in terms of each input feature’s contribution, commonly done by finding the output’s sensitivity to changes in input features, either through gradients or permutations [47, 39, 33, 37, 10].
 For images, attributions are visualized as pixel-level heatmaps.
@@
 Thus, proxies—such as checking if important pixels come from foreground regions or looking for significant model degradation after deleting important pixels—are used to study faithfulness.
-Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [Adebayo2018SanityCF, Frye2020ShapleyManifold, Sixt2019WhenEL, Hase2021TheOP].
-The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [Nguyen2021TheEO, colin2022what].
-Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [net2vec, netdiss, kimTCAV, zhou2014object].
-To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [Ramaswamy2022OverlookedFI].
+Despite strong performance on the proxies, clever sanity checks have revealed faithfulness problems [1, 15, 41, 22].
+The resulting heatmaps also lack interpretability—they only indicate where the model attends, not what semantics it recognizes [35, 7].
+Concept-based explanation methods (CBEMs) improve interpretability by explaining predictions in terms of human-understandable concepts (e.g., edges, colors, object parts) [14, 4, 24, 55].
+To recognize concepts, supervised CBEMs require a concept-annotated image dataset, which is costly and subjective; moreover, any fixed concept vocabulary may bias the explanation and miss important aspects of the model’s computation, raising faithfulness concerns [38].
 To avoid these issues, unsupervised CBEMs (U-CBEMs) discover concepts automatically as concept activation vectors (CAVs)—directions in the model’s intermediate representation space—paired with a concept importance score that captures the concept’s relevance to the output.
@@
 Motivated by Ribeiro et al.
-[lime], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$.
+[39], $E$ is completely faithful to $\phi$ if $s$ reproduces $\phi(\mathbf{X})\;\forall\;\mathbf{X}\in\mathcal{X}$.
 Thus, a faithfulness measure is defined by a choice of surrogate $s(.)$ and metric $d(.)$ as:
@@
 U-CBEMs: To generate explanations, U-CBEMs find, for all outputs $i\in C$, a set of $K$ CAVs $V_{i}=\{\mathbf{v}_{i,k}\}_{k=1}^{K}$ and concept importances $A_{i}=\{\alpha_{i,k}\}_{k=1}^{K}$.
-Note that while $K$ can vary per output, all prior works find the same $K$ per output.
+Note that while, for a given input, $K$ can vary per output class, all prior works find the same $K$ per output class.
 U-CBEMs also define a mechanism $\mathcal{P}(\mathbf{h};V):\mathcal{H}\to\mathbb{R}^{K}$ to project embedding h to the concept space defined by CAVs in $V$.
@@
 3.
-3 Related Works
-Unsupervised Concept-Based Explanation Methods.
+3 Related Works on U-CBEMs
 Wang et al.
-[Wang2015UnsupervisedLO] find CAVs by K-Means clustering intermediate embeddings computed on a dataset.
+[46] find CAVs by K-Means clustering intermediate embeddings computed on a dataset.
 Each cluster is associated with a semantic concept (e.g., car tire, window, headlight) by tracing back the clustered embeddings to their corresponding image patches.
 Ghorbani et al.
-[Ghorbani2019ACE] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate.
+[18] improve this with ACE, which finds concepts at different scales by hierarchically segmenting images into “candidate” segments and clustering the embeddings of each candidate.
 To ensure that the discovered CAVs fully represent the model’s behavior, Yeh et al.
-[ConceptSHAP] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [Young1985MonotonicSO] to find concept importance.
-ICE [Zhang2020InvertibleCE] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation.
+[48] propose ConceptSHAP (C-SHAP), which finds CAVs that maximize a proposed completeness score; they use Shapley Values [50] to find concept importance.
+ICE [53] extends ACE by replacing clustering with Non-negative Matrix Factorization (NMF), resulting in a parts-based explanation.
 The authors show that concepts found with NMF are more interpretable than those found by K-Means or Principal Component Analysis (PCA).
 Building on this, Fel et al.
-[Fel2022CRAFTCR] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance.
-CDISCO [graziani2023concept] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance.
-MCD [vielhaben2023multidimensional] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
-Finding issues with MCD’s interpretability, HU-MCD [hu-mcd] incorporates Segment Anything Model [SAM] to discover more interpretable concepts.
+[13] present CRAFT, which recursively applies NMF through the model to decompose concepts into sub-concepts, and they use sensitivity analysis for finding concept importance.
+CDISCO [20] uses the singular value decomposition to find CAVs and a gradient-based method for finding concept importance.
+MCD [44] extends the notion of CAVs to a multidimensional linear subspace, rather than a single vector, allowing it to describe more of the model’s behavior with fewer concepts.
+Finding issues with MCD’s interpretability, HU-MCD [21] incorporates Segment Anything Model [27] to discover more interpretable concepts.
+Recently, sparse autoencoders (SAEs) have emerged as a promising, scalable U-CBEM for large language models, and recent attempts have applied them to interpret vision models [19, 31, 12].
+CBEMs and U-CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model [28, 25, 51, 36].
 Table 1: Prior deletion-based proxies (top) & surrogate-based measures (bottom) are organized according to their differing factors: Deletion Space (DS), Deletion Method (DM), Metric, Surrogate, and Concept Projection (CP).
@@
 C-SHAP Normalized Top-1 Accuracy MLP Reconstruct $\to$ Original Model U-CBEM
-Concept Bottleneck Models.
-CBEMs should not be confused with Concept Bottleneck Models (CBMs), which create an inherently explainable concept model meant to replace the original black-box model.
-These methods [pmlr-cbm-koh20a, prob-cbm, posthoc-cbm, labelfree-cbm] train a model to first predict a set of interpretable concepts (i.e., a concept bottleneck); then, they use the concept representation for the task (e.g., object classification).
-By doing so, CBMs fully decompose the prediction into interpretable concepts.
-While CBMs are a promising approach, they typically fail to reach the same performance as black-box models, unless they introduce uninterpretable, “residual” concepts.
-As CBMs are inherently explainable, the notion of faithfulness does not extend to them.
-This work discusses faithfulness only for U-CBEMs.
 4 Unifying U-CBEM Faithfulness Measures
@@
 3.
+Appendix Sec.
+A offers a discussion and comparison with faithfulness metrics introduced for SAEs.
 4.1 Deletion-Based Proxies
@@
 This baseline must entirely delete the presence of the concept without affecting other concepts.
-However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [Frye2020ShapleyManifold].
+However, the feature attribution literature has shown that common baseline choices do not guarantee that a feature is completely deleted without affecting other features, leading to unfaithful explanations [15].
 Removing features by marginalization is more accurate but requires evaluating expensive, high-dimensional expectations.
 Second, there are no guarantees that representations in the Deletion Space will stay on manifold after deletion.
-The feature attribution literature [Frye2020ShapleyManifold, akata2022manifold, taufiq2023manifold, Slack2019FoolingLA, yeh2022threading] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold.
+The feature attribution literature [15, 2, 43, 42, 49] has found that the perturbed inputs used for creating explanations may be unrealistic, lying off the data manifold.
 In the off-manifold regions of the data space, a highly non-linear model behaves unpredictably, so the model’s outputs may not be meaningfully related to input features.
-Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [Frye2020ShapleyManifold].
-More seriously, these explanations can be adversarially attacked; in [Slack2019FoolingLA], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs.
-Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [Frye2020ShapleyManifold, taufiq2023manifold, vu2022emap, zaher2024manifold, Hase2021TheOP].
+Thus, calculating feature importance with off-manifold inputs can lead to incorrect explanations [15].
+More seriously, these explanations can be adversarially attacked; in [42], a model’s off-manifold behavior was modified to hide its dependence on undesirable features used for on-manifold inputs.
+Such issues can be mitigated by using on-manifold input perturbations, but this requires accurate modeling of the input distribution [15, 43, 45, 52, 22].
 These problems are not unique to feature attribution; they extend to deletions in the concept and weight space, which make similar manipulations in the model’s intermediate representation space.
@@
 4.2.2 Prior Surrogate-Based Measures
-Two prior works have introduced surrogate-based measures:
+Two prior works have proposed surrogate-based measures:
 ICE-Eval: ICE introduces its own evaluation measure (we term as ICE-Eval).
@@
 ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
-[…]
+SURF (Ours) 100 1.00 0.00 0.00 0.000 0 200
 Rand Imp C-SHAP-Eval (CEL) 9.02 -0.02 1.27 1.97 0.865 1M 205M
@@
 ICE-Eval 100 1.00 0.00 0.00 0.000 0 614K
-[…]
+SURF (Ours) 97.5 0.13 0.83 2.70 0.862 0 200
 Full Rand C-SHAP-Eval (CEL) 97.6 0.02 181.7 168.2 0.125 1M 205M
@@
 ICE-Eval 3.3 0.00 1.00 3.17 0.882 0 614K
-[…]
+SURF (Ours) 1.3 0.00 1.01 3.17 0.883 0 200
 5 Method: Surrogate Faithfulness (SURF)
@@
 If the model uses global pooling to reduce the final embedding of $\mathbf{x}$ to a single vector, the summation over $j$ above can be omitted.
-We define SURF’s surrogate $s$ to take the form of the linear layer (Eq. 9), where we use the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
+We define SURF’s surrogate $s$ to take the form of (Eq. 9), replacing $\mathbf{F}$ with the concept representation and $A_{i}$’s obtained from any U-CBEM. The surrogate is:
 $\mathbf{\hat{y}}_{i}$ $=s_{i}(.)=\sum_{j=1}^{HW}\sum_{k=1}^{K}\alpha_{i,k}\mathcal{P}(\mathbf{h}_{j};V_{i})_{k}\quad\forall\ i$ (10)
@@
 Following prior U-CBEMs and evaluation measures, SURF also operates only on the model’s final linear layer.
-Thus, its surrogate fully represents the model’s computation, while greatly reducing complexity compared to prior faithfulness surrogates.
+Thus, its surrogate fully represents the model’s computation, while not introducing any trainable parameters and greatly reducing complexity compared to prior faithfulness surrogates.
 5.2 The SURF Metrics
@@
 $\text{SURF}_{\text{EMD}}=\frac{1}{2|\mathcal{V}|}\sum_{\mathbf{x}\in\mathcal{V}}\sum_{i=1}^{C}|p_{i}-\hat{p}_{i}|$ (12)
-A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors. For regression, we only use $\text{SURF}_{\text{MAE}}$.
+We choose these specific metrics as they are easily interpretable.
+A faithful U-CBEM will have explanations with low $\text{SURF}_{\text{EMD}}$ and $\text{SURF}_{\text{MAE}}$ errors.
+For regression, we only use $\text{SURF}_{\text{MAE}}$.
 In contrast to prior surrogate-based measures, SURF meets all of our desiderata.
@@
 HU-MCD 1.97 0.384 99.7 0.149
+SAE 1.04 0.195 99.2 0.366
 (b) Multi-attribute prediction
@@
 2.83 96.6
+3.16 81.9
 (c) Age
@@
 32.6
+3.67
 6 Experiments
@@
 Then, we benchmark prior U-CBEMs across three tasks, revealing that SOTA methods are not faithful.
-Finally, we leverage SURF to analyze the trade-off between the number of concepts in an explanation and faithfulness.
+Finally, we leverage SURF to analyze the tradeoff between the number of concepts in an explanation and faithfulness.
 6.1 Measure-over-Measure Comparison
@@
 Then, we expect to obtain increasingly less faithful scores as we progressively increase randomness (from Rand Imp to Full Rand).
-We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [resnet] finetuned on the Caltech-101 dataset [caltech-101].
-Along with the SURF metrics, we report ICE’s Top-1 Accuracy (Top-1) defined in Eq.
-6 and ICE’s normalized L1 logit error (Norm L1) defined in Eq.
+We apply this sanity check on SURF, ICE-Eval, and C-SHAP Eval for an ImageNet-pretrained ResNet-50 [23] finetuned on the Caltech-101 dataset [30].
+Along with the SURF metrics, we report Top-1 Accuracy (Top-1) defined in Eq.
+6 and normalized L1 logit error (Norm L1) defined in Eq.
 5.
-As a complementary measure to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs.
-Rank Corr penalizes fine-grained changes that result in a different prediction ordering but ignores differences in magnitudes.
-Finally, we measure surrogate complexity by its learnable parameters and FLOPs needed to obtain the output from the concept representation.
-Since C-SHAP-Eval’s two-layer perceptron was originally trained with cross entropy loss (CEL), we also train the surrogate with L1 loss, as this may align better with the SURF metrics.
+As a complementary metric to $\text{SURF}_{\text{EMD}}$, we measure Spearman’s rank correlation (Rank Corr) between surrogate and model outputs.
+Rank Corr ignores magnitude differences but penalizes fine-grained changes that result in a different prediction ordering.
+Finally, we measure surrogate complexity by its learnable parameters and FLOPs from concept representation to output.
+Since C-SHAP-Eval’s surrogate was originally trained with cross entropy loss (CEL), we also train it with L1 loss, as this may align better with the SURF metrics.
 More details are in Appendix Sec.
-A.
+B.
 Results are reported in Tab.
@@
 In the Perfect setting, both ICE-Eval and SURF report perfect faithfulness, as expected.
-However, both C-SHAP-Eval variants are unable to achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations.
-In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; this occurs because their surrogates do not use concept importances, so using random concept importances has the same evaluation result as the Perfect setting.
+However, both C-SHAP-Eval variants do not achieve perfect faithfulness across any metric, showing C-SHAP-Eval’s limitations.
+In the Rand Imp setting, we observe that C-SHAP-Eval and ICE-Eval do not report differences when compared to the Perfect setting; because their surrogates do not use concept importances, using random concept importances do not change results from the Perfect setting.
 Contrast this with SURF, which clearly finds a deterioration in faithfulness across all metrics.
@@
 ICE-Eval and SURF report similarly poor faithfulness scores for this setting, as expected.
-Next, we turn our attention to the metrics.
-We highlight a failure case of Top-1 by looking at SURF’s results in the Rand Imp setting.
-The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the reconstructed output has severe errors on the remainder of the output distribution.
-Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the L1 error between logits.
+Next, we examine the metrics.
+SURF’s results in the Rand Imp setting highlight a failure case for Top-1.
+The Top-1 score is strong; however, the other metrics (especially $\text{SURF}_{\text{EMD}}$) reveal that the surrogate output has severe errors on the remainder of the output distribution.
+Norm L1 also comes with issues; notice how C-SHAP-Eval (CEL) reports a lower error compared to C-SHAP-Eval (L1) in the Perfect setting, despite C-SHAP-Eval (L1) directly minimizing the logit L1 error.
 We attribute this inconsistency to Norm L1’s emphasis on only the groundtruth class; while the L1 error across all classes (as measured by $\text{SURF}_{\text{MAE}}$) has decreased, the normalized error for the groundtruth class increased, giving an incomplete picture of faithfulness.
@@
 6.2 Benchmarking U-CBEMs
-We evaluate six prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD).
-We benchmark U-CBEMs on a varied range of tasks and models to demonstrate the applicability of SURF.
+We evaluate seven prior U-CBEMs (CDISCO, ICE, CRAFT, C-SHAP, MCD, HU-MCD, SAE) on a varied range of tasks and models to demonstrate the applicability of SURF.
 Specifically, our tasks are: 1) Object Classification, 2) Multi-Attribute Prediction, and 3) Age Regression.
@@
 CAVs and importances are found on the training set; the resulting explanations are evaluated on the test set.
-Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Sec.
-B.
+Dataset, finetuning, U-CBEM implementation details, and good-faith efforts made to adapt U-CBEMs to new tasks are in Appendix Sec.
+C.
 We evaluate on three tasks.
@@
 HU-MCD requires architectural modifications, so we can only evaluate it on this task, which uses a compatible ResNet-50.
-(2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [Sandler2018MobileNetV2IR] to predict the presence/absence of attributes on the CelebA dataset [celeba].
+Results on a VGG and InceptionV3 are in Appendix Sec.
+D.
+(2) Multi-Attribute Prediction: Given an image, we finetune a MobileNetV2 [40] to predict the presence/absence of attributes on the CelebA dataset [32].
 We report $\text{SURF}_{\text{MAE}}$ and attribute prediction accuracy (Attr-Acc), which is introduced as a task-specific replacement for Top-1.
-(3) Age Estimation: Given a human face, we finetune a ViT [VIT] to predict their age (in years) on the UTK-Face dataset [utkface].
+(3) Age Estimation: Given a human face, we finetune a ViT [9] to predict their age (in years) on the UTK-Face dataset [54].
 Because ViTs have negative activations, U-CBEMs with non-negativity assumptions (i.e., ICE and CRAFT) cannot be used.
@@
 In the evaluated settings, we observe that no prior U-CBEM is faithful to the original model.
-In the Object Classification task, the most faithful U-CBEM is HU-MCD with an $\text{SURF}_{\text{EMD}}$ of 0.384, denoting large errors in the probability-space.
-Only two of the tested U-CBEMs (ICE and MCD) perform significantly above random chance in the Multi-Attribute Prediction task.
-CDISCO exhibits an average error of 32.6 years in the Age Estimation task.
-We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP, all methods discover class-specific CAVs to reconstruct the embedding.
-Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an embedding from an image of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large logit- and probability-space errors.
+In the Object Classification task, the most faithful U-CBEM is SAE with an $\text{SURF}_{\text{EMD}}$ of 0.195, denoting significant errors in the probability-space.
+Only half of the tested U-CBEMs (ICE, MCD, and SAE) perform significantly above random chance in the Multi-Attribute Prediction task.
+SAE exhibits an average error of 3.67 years in the Age Estimation task.
+Note that this evaluation solely focuses on faithfulness and does not make any judgment on the interpretability side of the tradeoff.
+We believe that U-CBEMs are unfaithful for two reasons: (1) Other than C-SHAP and SAE, all methods discover class-specific CAVs to reconstruct the embedding.
+Crucially, the class-specific CAVs are found only on class-specific images, so using class $c_{1}$’s CAVs to reconstruct an image embedding of class $c_{2}$ will have large reconstruction error (e.g., the U-CBEM is operating out-of-distribution), and therefore, large SURF errors.
 This issue is exacerbated for CRAFT and ICE, which non-linearly project to the concept space.
 Though MCD and HU-MCD also find class-specific concepts, they achieve superior faithfulness by using concept subspaces; thus they faithfully explain more of the model with the same number of concepts.
-(2) Most methods do not satisfy a logit-based completeness criterion, preventing a simple reproduction of model outputs.
-Note that C-SHAP satisfies an accuracy-based completeness criterion and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be presented to the user in the explanation.
+(2) Most methods do not satisfy a logit-based completeness criterion, preventing a targeted reproduction of model outputs.
+C-SHAP, instead, satisfies an accuracy-based completeness criterion, and ICE, MCD, and HU-MCD satisfy the logit-based criterion by including uninterpretable residuals, which cannot be shown in the user explanation.
 Figure 2: We fit U-CBEMs on Sec.
@@
 However, explanations containing too many concepts may be difficult to interpret.
-Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [2010TheMM, Miller1956TheMN].
+Cognitive psychology studies show that humans can hold a limited number of items in their working memory at a time [8, 34].
 Thus, simple, or parsimonious, explanations are preferred, which communicate the bulk of the model’s computation with a few concepts.
@@
 SURF allows us to intelligently analyze this tradeoff.
-Using the Object Classification task from Sec.
-6.2, we fit U-CBEMs with an increasing number of concepts.
-Each U-CBEM is evaluated using the SURF measure, visualized in Fig.
+Using the Object Classification task (Sec.
+6.2), we fit U-CBEMs with an increasing number of concepts, each evaluated with the SURF measure and visualized in Fig.
 2.
-We observe that some U-CBEMs (CDISCO, CRAFT, C-SHAP) either do not improve or perform worse on one or both metrics as they discover more concepts.
-ICE marginally improves as we add more concepts.
+Some U-CBEMs (CDISCO, CRAFT, C-SHAP, ICE) either marginally improve or perform worse on one or both metrics as they discover more concepts.
+SAEs initially improve but then oscillates.
 Only MCD and HU-MCD uniformly improve as the number of concepts increases, and they exhibit a plateauing effect, marking a natural choice for the number of concepts.
-Interestingly, HU-MCD is more faithful than MCD only when both discover relatively few concepts; as the number of concepts grows, MCD becomes more faithful.
-This offers a more nuanced perspective on the findings from the HU-MCD paper, where they claimed superior faithfulness over MCD based on results from their deletion-based proxy.
+Interestingly, HU-MCD is more faithful than MCD only when both discover few concepts; as the number of concepts grows, MCD becomes more faithful.
+This adds a nuanced touch to the findings from the HU-MCD paper, which claimed superior faithfulness over MCD based on results from a deletion-based proxy.
 7 Conclusion
 This paper argues that we lack a clear view on how faithful current U-CBEMs really are, largely because of measures that deviate from the formal definition of faithfulness.
-In accordance with this definition, we develop desiderata for a faithfulness measure and propose SURF as a simple, principled measure for faithfulness.
-SURF passes a basic sanity check, reveals that state-of-the-art U-CBEMs are substantially unfaithful, and is a practical tool for studying the tradeoff between parsimony and faithfulness when choosing the number of concepts.
-We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores alongside their interpretability claims.
-[…]
+In accordance with this definition and associated desiderata, we propose SURF as a simple, principled measure for faithfulness.
+Among other things, we find that SOTA U-CBEMs do not faithfully explain the final output layer.
+While SURF accurately measures U-CBEM faithfulness, we emphasize that it should be paired with an interpretability analysis to judge the U-CBEM’s overall quality and value.
+We urge future work on concept-based explanations to adopt SURF as a standard faithfulness measure and to report SURF scores along their interpretability claims.
+U-CBEMs are predominantly applied on the final layer, but there is increasing interest in interpreting intermediate layers.
+SURF, as formulated, does not apply to this case, given the non-linear relationship between the explanation and the model’s output.
+Extending SURF to evaluate U-CBEMs for intermediate layers is important future work.
+8 Acknowledgment
+The support of the Office of Naval Research under grant N00014-24-1-2169, and IBM-Illinois Discovery Accelerator Institute (IIDAI), and USDA National Institute of Food and Agriculture under grant AFRI 2020-67021-32799/1024178 are gratefully acknowledged.
+This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No.
+DGE 21-46756.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
+Shubham Kumar also gratefully acknowledges the support of UIUC ECE’s Distinguished Research Fellowship and Promise of Excellence Fellowship.
 Supplementary Material
@@
   1. 1.
-    Sec. A: Implementation details for the measure-over-measure study (Sec. 6.1).
+    Sec. A: Discussion of related SAE faithfulness metrics.
   2. 2.
-    Sec. B: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
-A Implementation Details for Measure-over-Measure Comparison
+    Sec. B: Implementation details for the measure-over-measure study (Sec. 6.1).
+  3. 3.
+    Sec. C: Experimental details for the evaluation of U-CBEMs done in Sec. 6.2.
+  4. 4.
+    Sec. D: Results on additional models for the experiment done in Sec. 6.2.
+A Discussion and Comparison with SAE Faithfulness Metrics
+There has been a significant effort in recent years to explore using SAEs for explainability, particularly on large language models [5].
+SAEs are typically trained to minimize embedding reconstruction error (e.g., mean squared error—MSE), which is the primary metric used to evaluate their faithfulness.
+The intuition is that if an SAE can accurately reconstruct the intermediate embedding it is explaining, then downstream operations on the reconstructed embedding should also be faithfully preserved.
+While reconstruction error is useful, it is a proxy for faithfulness; there is no clear mapping from reconstruction to downstream prediction error, making it hard to interpret the significance of a given MSE error.
+To accommodate this, recent work [16, 29] introduces downstream losses on model outputs to measure faithfulness.
+Common choices are KL divergence and cross entropy loss.
+However, we argue that these are uninterpretable (as acknowledged in [16]).
+These metrics work well when used relatively, which fits the referenced papers’s need to compare different SAEs, but these metrics do not give an absolute measure of faithfulness (unlike SURF), which is necessary when one wants to make an independent assessment of a method’s faithfulness.
+Similarly, [31] develops a new SAE architecture (PatchSAE) and evaluates downstream classification accuracy after masking SAE latents (which is equivalent to a deletion-based approach).
+This metric suits their purpose, since they use SAEs to analyze model behavior, rather than to explain a prediction.
+[12] introduce a more robust SAE architecture (Archetypal SAEs) and introduce several metrics to evaluate.
+The closest metric to SURF is a plausibility metric, which measures the average maximum cosine similarity between the discovered CAVs and each classification vector.
+However, we argue that this metric does not give a holistic picture of faithfulness, since it only looks at the maximum and it is unclear how this metric maps on to errors in the output space (e.g., how does it get transformed by a softmax).
+We believe that SURF is a complementary measure that can be used to help study faithfulness of SAEs applied to vision models with the intention of explaining the model’s output computation.
+B Implementation Details for Measure-over-Measure Comparison
 This section contains details for the measure-over-measure comparison conducted in Sec. 6.1.
-A.1 Experimental Setting Details
+B.1 Experimental Setting Details
 In the Perfect setting, we set CAVs and importances according to the weights of the final classification layer.
@@
 For the random settings, we sample each dimension of the CAVs independently from a standard normal distribution, and then make the CAV unit norm. We sample concept importances uniformly from $[0,1)$.
-A.2 Surrogate Implementation
+B.2 Surrogate Implementation
 Recall that U-CBEMs define their own mechanism $\mathcal{P}:\mathcal{H}\to\mathbb{R}^{K}$ to project embeddings to the concept basis.
@@
 Recall in this case that $K=1$ and there is only one embedding (obtained from global pooling).
-A.2.1 SURF Surrogate (Ours)
+B.2.1 SURF Surrogate (Ours)
 Following Sec.
@@
 After the surrogate reconstructs the output, we add the constant bias term before evaluation.
-A.2.2 ICE Surrogate
+B.2.2 ICE Surrogate
 ICE’s surrogate takes the concept representation and reconstructs the embedding, as specified by the U-CBEM.
@@
 That is, $\psi(\mathbf{p_{i}})=\sum_{k=1}^{K}p_{i,k}\mathbf{v}_{i,k}$.
-A.2.3 C-SHAP Surrogate
+B.2.3 C-SHAP Surrogate
 C-SHAP’s surrogate defines a two-layer MLP $\psi_{\text{C-SHAP}}(.)$ to reconstruct the embedding, and then passes through the final classification layer to obtain outputs. The surrogate is:
@@
 Since C-SHAP’s surrogate contains learnable parameters.
-Following [ConceptSHAP], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities.
-We train the surrogate using the Adam optimizer [kingma2014adam] and a starting learning rate of $0.1$.
+Following [48], the MLP’s hidden layer has a dimensionality of 500, and the surrogate is trained using Cross Entropy Loss between the surrogate’s predicted class probabilities and the original model’s class probabilities.
+We train the surrogate using the Adam optimizer [26] and a starting learning rate of $0.1$.
 The surrogate is trained for at most 100 epochs, during which the learning rate is decayed by a factor of 0.5 when the training loss plateaus.
 We stop training when the learning rate is reduced below $10^{-7}$.
-A.3 FLOPs Computation
+B.3 FLOPs Computation
 We treat multiply and add as 2 FLOPs. Let $K$ be the number of concepts, $C$ be the number of output classes, and $D$ by the dimensionality of the embedding space.
@@
 Thus, the number of FLOPs required is $2C(H_{1}K+H_{1}D+D)$, where $H_{1}$ is the dimensionality of the MLP’s hidden layer.
-B Additional Details for Benchmarking U-CBEM Faithfulness
+C Additional Details for Benchmarking U-CBEM Faithfulness
 Sec. 6.2 evaluates prior U-CBEMs, finding that they produce unfaithful explanations. This section gives experimental and implementation details for this evaluation.
-B.1 Model Details
+C.1 Model Details
 Object Classification: Our evaluation uses a ResNet-50, available from the timm library.
@@
 The model is trained with data augmentation and using L1 Loss.
-B.2 Dataset Details
+C.2 Dataset Details
 We learn each class’s concepts and importances using images of the class from the training set.
@@
 As UTK-Face does not have any classes, all U-CBEMs find CAVs on all embeddings.
-B.3 U-CBEM Details
+C.3 U-CBEM Details
 This section specifies implementation details for each U-CBEM that was benchmarked.
-B.4 CDISCO
+C.4 CDISCO
 The open-source implementation provided by the authors was used to evaluate this approach.
@@
 According to the paper and the code used for multi-class classification, we re-implemented CDISCO to work in the case of single-output regression.
-B.5 ICE
+C.5 ICE
 The open-source implementation provided by the authors was used to evaluate this approach.
@@
 Due to the potentially large number of embeddings, we use the batched-version of NMF.
-Concept importance is found using the TCAV method proposed in [kimTCAV].
+Concept importance is found using the TCAV method proposed in [24].
 Embeddings are projected to the concept space using sci-kit learn’s transform method.
@@
 Thus, CAVs and concept importances were learned for each task independently.
-B.6 CRAFT
-The open-source implementation provided in the Xplique toolbox [fel2022xplique] (by the same authors) was used to evaluate this approach.
+C.6 CRAFT
+The open-source implementation provided in the Xplique toolbox [11] (by the same authors) was used to evaluate this approach.
 CRAFT uses NMF to find CAVs and introduces a Sobol-based sensitivity approach to find concept importance.
 Embeddings are projected to the concept space using the toolbox’s built-in transform method.
-B.7 ConceptSHAP
+C.7 ConceptSHAP
 ConceptSHAP differs from previous U-CBEMs because it discovers CAVs globally (i.e., class-agnostic).
@@
   2. 2.
-    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([shapley]) and is class-specific.
+    Added more efficient concept importance scoring via Shapley Values. This is implemented using KernelSHAP ([33]) and is class-specific.
   3. 3.
     Fixed the implementation of the completeness score
-B.8 MCD
+Table 5: Additional Benchmark Results.
+We apply SURF to evaluate explanations from prior U-CBEMs on the Object Classification task from Sec.
+6.2 on a VGG11 and InceptionV3.
+Along with the $\text{SURF}_{\text{MAE}}$ and $\text{SURF}_{\text{EMD}}$, we report other metrics to serve as a comparison.
+Our trends hold: prior U-CBEMs are not faithful, as indicated by large errors in the logit and probability space.
+(a) Object classification
+U-CBEM $\text{{SURF}}_{\text{MAE}}$ ($\downarrow$) $\text{{SURF}}_{\text{EMD}}$ ($\downarrow$) Top-1 (%) ($\uparrow$) Rank Corr ($\uparrow$)
+CDISCO 2.89 0.936 0.4 -0.01
+ICE 3.58 0.616 93.7 0.498
+CRAFT 2.50 0.844 51.0 0.214
+C-SHAP 3.28 0.884 1.1 -0.00
+MCD 1.95 0.738 83.3 0.306
+HU-MCD
+SAE 0.47 0.567 62.4 0.504
+(b) Object classification
+$\text{{SURF}}_{\text{MAE}}$ $\text{{SURF}}_{\text{EMD}}$ Top-1 (%) Rank Corr
+4.08 0.847 1.52 0.012
+3.34 0.595 50.3 -0.05
+4.26 0.652 8.41 0.131
+4.31 0.754 9.70 -0.08
+2.76 0.327 89.1 0.385
+0.35 0.226 72.0 0.770
+C.8 MCD
 Instead of representing concepts as single CAVs, MCD represents concepts as multi-dimensional subspaces.
-In [vielhaben2023multidimensional], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace.
+In [44], any concept $l$ for class $i$ is characterized by a representative basis $C^{i,l}=\{\mathbf{v}_{j}^{i,l}|j=1...d_{l}\}$, where $d_{l}$ denotes the dimensionality of the subspace.
 To ensure the union of all concepts span the entire feature space, they define the orthogonal complement concept $C^{i,\perp}=\text{span}(C^{i,1},...,C^{i,K})^{\perp}$, where $K$ is the number of concepts originally discovered, letting $C^{i,K+1}\equiv C^{i,\perp}$.
@@
 We use the open-source implementation provided by the authors.
-B.9 HU-MCD
+C.9 HU-MCD
 HU-MCD directly extends upon MCD, adopting the same framework of concept subspace, orthogonal complement, and unique decomposition onto the concept basis.
@@
 We use the open-source implementation provided by the authors.
+C.10 SAE
+We learn a Top-K SAE [16] on the training embeddings (before average pooling) using the Overcomplete library.
+SAEs learn class-agnostic concepts (in contrast to U-CBEMs that learn CAVs for a specific class output); to enable a fair comparison, our SAE learns the same number of total CAVs as the other U-CBEMs for each setting.
+As an example, our SAE for the Object Classification task learns 505 CAVs (5 per class output, 101 total classes).
+At reconstruction time, we set the $K$ in the Top-K SAE to 5, directly mirroring the other U-CBEMs.
+D Additional Benchmark Results
+We repeat our experiment on the Object Classification task from Sec.
+6.2 on a VGG11 and InceptionV3 to confirm trends hold.
+We use the ImageNet pre-trained weights, and finetune them for the Object Classification task.
+Results are reported in Tab.
+5.
+For both models, SAE is the most faithful.
+For VGG11, we find that all U-CBEMs tend to exhibit large EMD errors (larger than 0.5).
+Though SAE is able to reduce the logit error significantly, the metrics indicate that there are still issues when trying to reproduce predictions in the probability space.
+While ICE can accurately reconstruct the Top-1 prediction, it has large logit errors.
+For InceptionV3, we find ICE is not faithful, and instead MCD is second-best (after SAE).
+Compared to the ResNet studied in Sec.
+6.2, we believe that the models tested here have more nuance in their predicted probability space, which may lead to larger EMD errors, even if the logit error seems small.